### v1

# Single-Channel EEG Tokenization Through Time-Frequency Modeling

Abstract.
We introduce TFM-Tokenizer, a novel tokenization framework tailored for EEG analysis that transforms continuous, noisy brain signals into a sequence of discrete, well-represented tokens for various EEG tasks. Conventional approaches typically rely on continuous embeddings and inter-channel dependencies, which are limited in capturing inherent EEG features such as temporally unpredictable patterns and diverse oscillatory waveforms. In contrast, we hypothesize that critical time-frequency features can be effectively captured from a single channel. By learning tokens that encapsulate these intrinsic patterns within a single channel, our approach yields a scalable tokenizer adaptable across diverse EEG settings. We integrate the TFM-Tokenizer with a transformer-based TFM-Encoder, leveraging established pretraining techniques from natural language processing—such as masked token prediction—followed by downstream fine-tuning for various EEG tasks. Experiments across four EEG datasets show that TFM-Token outperforms state-of-the-art methods. On TUEV, our approach improves balanced accuracy and Cohen’s Kappa by $5\%$ over baselines. Comprehensive analysis of the learned tokens demonstrates their ability to capture class-distinctive features, enhance frequency representation, and ability to encode time-frequency motifs into distinct tokens, improving interpretability.
EEG, Tokenization, Deep learning
journalyear: 2025copyright: acmcopyrightconference: 2025; ; ccs: Computing methodologies Machine learning approachesccs: Applied computing Health care information systems
1. Introduction
Electroencephalograms (EEGs) reveal real-time neuronal activity with millisecond precision, reflecting the responses to various event stimuli. This makes EEGs essential not only for fundamental research (Conrad et al., 2019; Li et al., 2021) but also for diverse applications such as sleep staging (Phan et al., 2022; Yang et al., 2023; Pradeepkumar et al., 2024), neurological disease detection (Kotoge et al., 2024; Afzal et al., 2024), and brain-computer interfaces (Precup and Bachman, 2012; Wolf et al., 2022; Li et al., 2023). Practical EEG analysis is a complex, multi-step process involving filtering, feature engineering, pattern recognition, annotation, etc. Deep learning (DL) models have shown remarkable success in automating EEG analysis across various tasks (Sharma et al., 2022; Tang et al., 2022; Chen et al., 2023). These achievements stem from their representation learning capability, projecting noisy EEGs onto a discriminative feature space that captures associations with discrete neurophysiological events. Recent neuroscientific literature reports that DL methods successfully match the reliability and accuracy of 20 clinical experts in identifying seizure-related events (Jing et al., 2023).
Figure 1. TFM-Tokenizer tokenizes EEG signals by encoding time-frequency motifs as discrete tokens. The neural tokenizer used for LaBraM (Jiang et al., 2024), which is a raw signal-based tokenizer, fails to capture time-frequency variations. Our TFM-Tokenizer effectively preserves this information.
Despite their success, effectively representing EEGs remains a primary focus in the research community. Real-world EEG recordings are acquired under diverse scenarios, using different devices, and even for the same task, data formats often vary due to mismatched channels and variable lengths (Yang et al., 2024). Unfortunately, most existing methods typically learn representations on a case-by-case basis with specific architectures or fixed channel settings. These methods exhibit limited generalization across tasks and poor scalability to different data formats. For example, a sleep apnea model is not readily applicable to a sleep staging task, since the former aims to capture anomaly patterns while the latter is to learn discriminative representations (Pradeepkumar et al., 2024). Likewise, a sleep staging model trained on central region EEGs cannot directly generalize to frontal EEG data. As a result, retraining, elaborate fine-tuning, or model adjustments are often required (Yang et al., 2024), escalating practical costs and further restricting applicability to advanced devices or non-standard data, such as ECoG (Pradeepkumar et al., 2021). There is thus an urgent need to develop an EEG analysis method that serves broader research objectives.
Recently, the transformative impact of large foundation models(Touvron et al., 2023; Achiam et al., 2023) has inspired and elevated EEG representation learning to new heights. Several foundation EEG models have been proposed (Duan et al., 2023; Yi et al., 2024; Mohammadi Foumani et al., 2024; Yang et al., 2024), demonstrating both enhanced performance and generalization. Researchers often tokenize EEGs into discrete, short-duration snapshots across different data formats and model their dependencies leveraging foundational yet powerful Transformers. By leveraging the paradigm of pretraining on large-scale, cross-task datasets, these models can learn universal representation, enabling effective adaptation to downstream tasks without requiring task-specific architectures (Kotoge et al., 2024; Mohammadi Foumani et al., 2024). Thanks to unified tokenization and the automated input adaptation of the Transformer, such methods also address challenges such as complex data format reconfigurations (Yang et al., 2024) and labeled data scarcity (Jiang et al., 2024). However, this direction remains nascent, and several limitations remain:
  • Inappropriate Tokenization Representation. One reason large language models (LLMs) succeed is their effective tokenization and similar benefits have been demonstrated in image (Van Den Oord et al., 2017) and video (Wang et al., 2024b; Agarwal et al., 2025) tokenization. However, existing foundation EEG models generally do not adopt a discrete tokenization paradigm. Although some methods claim to provide an EEG “tokenizer,” they typically lack a discrete approach similar to NLP. For instance, LaBraM (Jiang et al., 2024) employs a neural tokenizer only during pretraining but relies on raw EEG signals during inference. BIOT (Yang et al., 2024) suggests a “biosignal tokenizer” that remains continuous rather than discretized. Although time-series tokenization methods have shown promise (Ansari et al., 2024; Rasul et al., 2023), they do not scale well to EEG’s higher sampling rates and other artifacts.
  • Insufficient Frequency Representation. Capturing eventful EEG features, which are characterized by distinct frequencies, is a primary focus of EEG analysis. However, tokenizing and modeling raw EEGs often lead to a loss of frequency diversity. This frequency representation collapse is a common issue in time-series modeling, as low-frequency components typically dominate the EEG data, biasing models toward lower frequencies while overlooking critical high-frequency features, such as spikes. Although LaBraM has recently incorporated frequency reconstruction into its objective function, it still relies on raw EEG signals, inevitably resulting in sub-optimal representation, as shown in Figure 1.
  • Scalability and Generalization. EEG-related tasks vary significantly in channel configurations. For example, seizure detection typically utilizes 16 channels, whereas sleep studies often require only 1–2 channels. However, existing models are primarily designed for multi-channel settings, heavily relying on cross-channel prediction. This design limits their scalability and adaptability to configurations with fewer or even single channels, as well as to varying acquisition setups.
Therefore, in this paper, we propose TFM-Token, an effective, fully discretized EEG tokenization framework that captures time-frequency motifs from single-channel EEG signals into distinct tokens. Technically, our contributions are as follows:
  • TFM-Tokenizer and TFM-Encoder: We introduce a scalable discrete tokenization framework for EEG, transforming single-channel signals into discrete token sequences akin to NLP models. TFM-Tokenizer converts EEG into discrete tokens, and TFM-Encoder utilizes them for downstream tasks.
  • Joint Modeling of Frequency and Temporal Dynamics: Our tokenizer integrates raw EEG patches with time-frequency representations, using frequency band and temporal masking to capture essential frequency patterns while disentangling temporal variations (Figure 1).
  • Scable tokenization: Our single-channel approach enables flexible adaptation across EEG tasks and channel configurations. TFM-Tokenizer further enhances state-of-the-art EEG models, such as LaBraM (Jiang et al., 2024).
  • Empirical Validation and Token Quality Analysis: We evaluate our framework on four EEG downstream tasks, demonstrating state-of-the-art performance. Beyond performance, we comprehensively analyze token quality, including token visualization, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
2. Related work
EEG Representation Learning. To learn general representations and address issues of label scarcity in EEG data, self-supervised learning (SSL) has emerged as a prominent paradigm, and existing works can be categorized into two main approaches: contrastive learning and self-prediction. Contrastive learning methods, including TS-TCC (Eldele et al., 2021), TF-C (Zhang et al., 2022) and SplitSEE (Kotoge et al., 2024), leverage augmentation or transformation of EEG inputs to learn consistent representations. In contrast, self-prediction methods, namely BENDR (Kostas et al., 2021), MAEEG (Chien et al., 2022), BIOT (Yang et al., 2024) and EEG2REP (Mohammadi Foumani et al., 2024), aim to accurately reconstruct masked or corrupted input. However, their learning objectives heavily rely on cross-channel prediction to focus on spatial characteristics. In contrast, our method emphasizes inherent time-frequency features within a single-channel setting and can adapt to any channel configuration.
Foundation EEG Models. Inspired by the success of foundation models in NLP, recent efforts have sought to develop foundation models for EEG analysis. These models can be categorized into decoding and encoder-based methods. Decoding-only methods focus on generative tasks like EEG-to-text translation, with representative works including DeWave (Duan et al., 2023), EEG2Text (Liu et al., 2024), and E2T-PTR (Wang et al., 2024a). In contrast, encoder-only methods concentrate on fundamental EEG classification tasks and representation learning. Notable models include LaBraM (Jiang et al., 2024), BIOT (Yang et al., 2024), BRANT (Zhang et al., 2024), and MMM (Yi et al., 2024). Our work aligns with this latter category, focusing on enhancing the representation quality to improve classification performance.
EEG Tokenization. Tokenization has been instrumental in NLP, where discrete subword units have proven to reduce data complexity and improve model performance and interoperability. Existing attempts for EEGs include patch-based continuous tokenization, such as BIOT (Yang et al., 2024) and BRANT (Zhang et al., 2024), and vector quantization (VQ)-based methods like DeWave (Duan et al., 2023). Patch-based methods do not involve encoding or quantization, leading to unbounded and continuous representations that lack distinctiveness and interpretability. In contrast, VQ-based tokenizers, traditionally successful in tokenizing continuous images (Esser et al., 2020), have recently been adapted for EEG by LaBraM (Jiang et al., 2024), However, its tokenizer is utilized only during training and not during inference. Conceptually, its primary role is to pre-train classification layers, rather than encoding inputs and reducing data complexity. Here, our method is explicitly VQ-based, treating the codebook as a real tokenizer for EEG data. Moreover, we enforce each token to capture time-frequency motifs(Xu et al., 2023) in EEG inputs, ensuring a more structured and interpretable representation.
Frequency Representation Collapse. Frequency domain analysis is crucial in EEG and general time series analysis (Elvander and Jakobsson, 2020; Wu et al., 2021, 2023; Woo et al., 2022). In real-world signals, time-domain observations inherently mix multiple frequency components, and high-energy, low-frequency signals often dominate the spectrum (Huang Norden E Shen Zheng and H, 1998; Lai et al., 2018). As a result, these entangled frequency features makes it difficult for models to distinguish between them (Zhou et al., 2022; Piao et al., 2024). Recent studies have shown that these entangled signals can lead to a collapse in the learned frequency representations (Zhi-Qin John Xu et al., 2020; Piao et al., 2024). Models tend to overemphasize the dominant low-frequency features while neglecting the high-frequency details. This issue can lead to a lack of capturing various EEG waveforms and degenerating data representation (Park and Kim, 2022). Motivated by these works, our paper focuses on developing methods to learn diverse, informative frequency features. In Section 3, we provide an analysis of our proposed frequency-domain tokenizer and its impact on model performance.
3. Preliminary
Figure 2. Overview of the TFM-Token framework. (a) TFM-Tokenizer Pretraining: Through dual-path encoding and masked prediction, learns to capture time-frequency motifs into discrete tokens. (b) TFM-Encoder Pretraining: Trains on learned EEG tokens using masked token prediction. (c) Masking Strategy: A combination of frequency band masking and temporal masking is used for TFM-Tokenizer pretraining. (d) Localized Spectral Window Encoder: Processes individual spectral windows from $\mathbf{S}$, extracts frequency band information, and aggregates features across all bands into a single compact embedding per window.
3.1. Notations and Problem Statements
EEG Data. Let $\mathbf{X}\in\mathbb{R}^{C\times T}$ denote a multi-channel EEG recording with $C$ channels and $T$ time samples. For each channel $c\in\{1,...,C\}$, we denote a single channel EEG by $x^{c}\in\mathbb{R}^{T}$. To capture complementary structures in both the time and frequency domains, we decompose $x^{c}$ into: (1) time-frequency representation $\mathbf{S}$ and (2) raw EEG patches $\{x_{i}\}_{i=1}^{N}$. For simplicity, we omit the channel index; henceforth, $x$ will refer to a single-channel EEG signal unless otherwise specified.
Short-Time Fourier Transform (STFT). To obtain the time-frequency representation, i.e.g, spectrogram, $\mathbf{S}$, we apply a STFT to $x$ using a windowing function $w(.)$ of length $L$ and a hop size $H$:
(1) $\mathbf{S}(\omega,\tau)=\left|\sum_{l=0}^{L-1}x(\tau H+l)w(l)e^{\frac{-j2\pi\omega l}{L}}\right|$
where $\omega$ indexes the discrete frequencies and $\tau$ indexes the time segments (i.e., time windows shifted by $H$). We retain only the magnitude |.||.|| . | to form $\mathbf{S}\in\mathbb{R}^{F\times N}$, where $F$ is the number of frequency bins and $N$ is the number of time windows.
Problem Statement 1 (EEG Tokenization): Given a single channel EEG $x$, we aim to learn a tokenization function
$f_{\text{tokenizer}}:\mathbb{R}^{T}\rightarrow\mathcal{V}^{N\times D}$
where $\mathcal{V}$ is a finite EEG token vocabulary of size $k$, and $D$ is the dimension of each token embedding. The tokenizer function $f_{\text{tokenizer}}$ should project $x$ (or transformations) into a sequence of discrete tokens $\{v_{i}\}_{i=1}^{N}$, where each $v_{i}\in\mathcal{V}$. These tokens represent various temporal and frequency “motifs”: meaningful EEG patterns characterized by distinct temporal and frequency characteristics. Therefore, $\mathcal{V}$ is learnbale from $\mathbf{S}$ and the temporal patches $\{x_{i}\}_{i=1}^{N}$.
Remark. We here hold several expectations for the learned motif tokens. First, these tokens are expected to reduce redundancy, noise, and complexity, providing a compact, sparse, and informative representation of EEGs. Second, these motifs should effectively capture essential neurophysiological patterns from both temporal and frequency domains. Third, the tokens should generalize well across different EEG tasks, enhancing the efficiency and interpretability of the data. We set up related research questions to evaluate these expectations in Section. 5.
Problem Statement 2 (Multi-Channel EEG Classification): Given EEGs $\mathbf{X}$ and a fixed, learned single-channel tokenizer $f_{\text{tokenizer}}$, we apply $f_{\text{tokenizer}}$ independently to each channel $c$ to obtain a tokenization representation $\Bigl{\{}\{v_{i}^{c}\}_{i=1}^{N}\Bigr{\}}_{c=1}^{C}$. Then, these tokens can serve various downstream tasks. For classification tasks, they are mapped to labels by:
$f_{\text{classifier}}:(\mathcal{V}^{D})^{N\times C}\rightarrow\mathbf{Y}$
where $Y$ is the target labels (e.g., EEG events, seizure types). By aggregating and processing the tokens across all channels, $f_{\text{classifier}}$ predicts a label $y\in\mathbf{Y}$. Notably, $f_{\text{classifier}}$ can be any downstream-oriented model, and its training is performed separately from the EEG tokenizer $f_{\text{tokenizer}}$.
4. Methodology
TFM-Token comprises two components: (1) TFM-Tokenizer: addresses problem statement 1 by converting continuous EEG signals into discrete tokens, capturing key time-frequency features, and (2) TFM-Encoder: tackles problem statement 2 by leveraging these tokens for downstream EEG tasks, such as classification. To mitigate the quadratic complexity of standard Transformers (Vaswani, 2017), we employ a linear attention mechanism (Katharopoulos et al., 2020; Wang et al., 2020) across all transformer modules. Given input EEGs, TFM-Tokenizer independently tokenizes each channel into discrete token sequences. The TFM-Encoder then processes these tokens to perform classification. For training TFM-Token, we first conduct an unsupervised pretraining of TFM-Tokenizer in a single-channel setting to encode time-frequency motifs into discrete tokens (see Figure 2a, Sec 4.1). The tokenizer is then frozen, and TFM-Encoder undergoes masked token prediction pretraining (see Figure 2b, Sec 4.2), followed by fine-tuning for downstream EEG tasks.
Time-Frequency discretization. The input to TFM-Token consists of discrete time and frequency patches, applied to both raw EEG data and its corresponding spectrogram. This approach addresses the transient nature of EEG waveforms, which may decompose the entangled patterns and is foundational for motif learning. Specifically, we segment $x$ into time-domain length patches $L$ with the same hop size of $H$. This yields a sequence of patches $\{x_{i}\}_{i=1}^{N}$, where $N=\lfloor(T-L)/H\rfloor+1$ is the total number of patches. Using the same $L$ and $H$ for both STFT and temporal segmentation, we ensure that each patch $x_{i}$ corresponds one-to-one with a spectral window $\mathbf{S}_{i}$. In practice, we obtain the raw EEG patches by configuring the convolutional layers with a kernel size and stride that match the STFT window length $L$ and hop size $H$. Additionally, we denote embeddings and masks as $\mathbf{E}$ and $\mathbf{M}$, respectively.
4.1. Single Channel TFM-Tokenizer
We introduce the TFM-Tokenizer, a scalable module for tokenizing single-channel EEG signals $x$ by effectively capturing both their temporal and frequency dynamics. Our design is inspired by the Vector-Quantized Variational Autoencoder (VQ-VAE)(Van Den Oord et al., 2017), which has been widely adopted for tokenization efforts in other domains such as video processing (Agarwal et al., 2025). As illustrated in Figure 2a, the TFM-Tokenizer comprises three primary components: (1) Localized Spectral Window Encoder, (2) Temporal Encoder, and (3) Temporal Transformer. At a high level, TFM-Tokenizer adopts a frequency-then-time paradigm. First, we isolate each spectral window $\mathbf{S}_{i}$ (for $i=1,…,N$) to capture intra-frequency band dependencies, yielding an aggregated representation for each window (Localized Spectral Window Encoder, see Figure 2d). Next, these frequency-based representations are combined with features extracted from the corresponding raw EEG patches $\{x_{i}\}_{i=1}^{N}$ (Temporal Encoder). Finally, the combined features are passed through the Temporal Transformer to capture long-range temporal dependencies, and the resulting embeddings are quantized into discrete tokens.
Localized Spectral Window Encoder. Capturing frequency-band characteristics or compositions is crucial for EEG analysis, as the data often contains oscillatory components (e.g., alpha, beta) with varying amplitudes and time. Unlike previous studies that project an entire spectral window with a single linear layer (Yang et al., 2024), we propose to patch along the frequency axis, enabling better modeling of cross-frequency dependencies. In detail, this encoder consists of three main steps (see Figure 2d):
  • Frequency Patch Encoder. Given a set of spectral windows $\{\mathbf{S}_{i}\}_{i=1}^{N}$, we isolate and divide each spectral window $\mathbf{S}_{i}$ into $P$ non-overlapping patches $\{\mathbf{S}_{(i,p)}\}_{p=1}^{P}$, each spanning $\Delta f$ frequency bins such that $P.\Delta f=F$. We then project each frequency patch into a latent space:
    $e_{(i,p)}=\text{GroupNorm}\left(\text{GeLU}\left(\mathbf{W}_{p}\mathbf{S}_{(i,p)}\right)\right)$
    where $\mathbf{W}_{p}\in\mathbb{R}^{D\times\Delta f}$ is a learnable matrix.
  • Frequency Transformer. In order to capture the interactions among the frequency patch embeddings $e_{(i,p)}$, we apply a frequency transformer that operates along the frequency axis within each spectral window $\mathbf{S}_{i}$. This enables the capturing of intra-frequency band dependencies (e.g., interactions between low and high-frequency bands).
  • Gated Patchwise Aggregation: In many EEG scenarios, large portions of the frequency spectrum can be irrelevant. For instance, tasks related to sleep primarily focus on frequency bands up to approximately 32 Hz (Chen et al., 2023). Also, the frequencies of interest vary across conditions and tasks. To emphasize important frequency patches and suppress the rest, we adopt a gated aggregation mechanism. Formally, we first group frequency patch embeddings, then apply gated weighting before concatenating them into a single embedding $\mathbf{E}^{F}_{i}$
    $\mathbf{E}^{F}_{i}=\text{Concat}\left[\sigma\left(\mathbf{W_{g1}e_{(i,p)}}\right)\mathbf{W_{g2}e_{(i,p)}}\right]$
    where $\mathbf{W_{g1}},\mathbf{W_{g2}}$ are trainable parameters and $\sigma(\cdot)$ is the element-wise sigmoid function.
Temporal Encoder. To capture temporal dynamics from the raw EEG patches $\{x_{i}\}_{i=1}^{N}$, we perform a linear projection of the patch followed by GELU(Hendrycks and Gimpel, 2016) activation and group normalization, producing temporal embeddings $\{\mathbf{E}^{T}_{i}\}_{i=1}^{N}$.
Temporal Transformer. We form a unified time-frequency embedding by combining each aggregated frequency embedding $\mathbf{E}_{i}^{F}$ with its corresponding temporal embedding $\mathbf{E}_{i}^{T}$ and feed the resulting vectors into a temporal transformer. This module attends jointly to time and frequency features across all windows, capturing long-range dependencies. The output is then quantized into discrete tokens $\{v_{i}\}_{i=1}^{N}$ using a learnable codebook $\mathcal{V}^{k}$. Notably, we do not add any positional encoding as EEG signals are inherently non-stationary and often exhibit chaotic behavior; thus, our goal is to capture their distinctive features without enforcing positional constraints. We further provide an ablation study (Appendix A.3) showing the effect of omitting positional encodings.
Tokenizer Codebook. Our tokenizer is designed to capture temporal frequency motifs by applying vector quantization along the time axis, that is, individual patches. This is different from conventional quantization methods in computer vision, which typically operate on the embedding dimension of a patch (Van Den Oord et al., 2017; Esser et al., 2020). As a result, each token in our vocabulary directly represents a waveform within a specific short duration and can facilitate the retrieval of the timestamp of input EEGs to understand the composition of an EEG. This enhances interpretability; for instance, specific codes in the vocabulary are activated in response to the occurrence of particular temporal or frequency patterns. To validate this, we conducted a visual inspection study, with results presented in Section 5.5.
Frequency Making Prediction for Tokenizer Learning. The input to the tokenizer is the full frequency band of a temporal patch. To facilitate effective frequency learning, we employ a frequency-band and temporal masking strategy during TFM-Tokenizer training. We partition $\mathbf{S}$ into $N_{F}=\lfloor\frac{F}{\delta_{f}}\rfloor$ frequency groups of size $\delta_{f}$ and randomly mask a subset based on a predefined mask ratio (e.g., 0.5), forming a frequency-band mask $M_{F}$. Simultaneously, a temporal mask $M_{T}$ is applied along the time axis. The final mask, $M=M_{F}\land M_{T}$, is used to mask the STFT representation, $S^{M}=M\odot S$, where $\odot$ denotes element-wise multiplication. For data augmentation and efficient training, we adopt a symmetric masking approach from (Jiang et al., 2024) by defining $M_{\text{sym}}=1-M$. Given $\mathbf{S}^{M}$ and $x$, our TFM-Tokenizer first quantized them into a sequence of distinct tokens, which is then passed through a transformer decoder followed by a linear decoder that predicts the masked regions. The tokenizer is trained using a masked reconstruction loss:
$\mathcal{L}_{\mathrm{rec}}=\sum_{(f,t)\in\mathcal{I}_{\mathrm{mask}}}\!\bigl{\|}\mathbf{S}(f,t)-\hat{\mathbf{S}}(f,t)\bigr{\|}_{2}^{2}$
where $(f,t)\in\mathcal{I}_{\text{mask}}$ represents masked frequency–time indices and $\hat{\mathbf{S}}$ is the reconstruction of $\mathbf{S}$. Alongside the masked reconstruction loss, we incorporate the codebook and commitment losses (Van Den Oord et al., 2017), yielding the total loss:
$\mathcal{L}_{\mathrm{token}}=\mathcal{L}_{\mathrm{recon}}+\alpha\;\sum_{i}\bigl{\|}\mathrm{sg}[E_{i}]\;-\;v_{i}\bigr{\|}_{2}^{2}+\beta\;\sum_{i}\bigl{\|}E_{i}\;-\;\mathrm{sg}[v_{i}]\bigr{\|}_{2}^{2}$
where, $\mathrm{sg}[\cdot]$ is the stop-gradient operator, and $\alpha,\beta$ are hyperparameters. We also employ exponential moving average strategy(Van Den Oord et al., 2017) for stable codebook updates.
4.2. Token-Wise TFM-Encoder
The TFM-Encoder is designed to aggregate tokenized representations across channels and perform downstream classification. It consists of two main components: (1) a token-embedding lookup table and (2) linear attention transformer layers. Given a multi-channel recording $\mathbf{X}\in\mathbb{R}^{C\times T}$, we apply the fixed (pretrained) TFM-Tokenizer to each channel independently, obtaining discrete token sequences $\Bigl{\{}\{v_{i}^{c}\}_{i=1}^{N}\Bigr{\}}_{c=1}^{C}$. Each token is mapped to its embedding via a lookup table initialized with the TFM-Tokenizer ’s learned codebook weights. We flatten the embeddings across channels and incorporate channel and position embeddings to encode respective information. We then prepend a [CLS] token (Devlin, 2018), the output of which is designed to capture the overall representation of the sequence for downstream tasks. The resulting embeddings are then processed by linear attention transformers in TFM-Encoder.
Token Prediction for TFM-Encoder Learning. We adopt a strategy akin to masked language modeling. We randomly mask tokens across multiple channels and time steps and train the model to predict these masked tokens via a cross-entropy loss. Along with representation learning, this approach enhances robustness to missing or corrupted data, common in real-world EEG systems where channels or time segments may be dropped or noisy. Finally, the TFM-Encoder is finetuned to downstream tasks.
5. Experiments and Results
Table 1. EEG classification performance comparison on TUEV and TUAB datasets.
Models Number TUEV (event type classification) TUAB (abnormal detection)
of Params Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC
SPaRCNet(Jing et al., 2023) 0.79M $0.4161\pm 0.0262$ $0.4233\pm 0.0181$ $0.7024\pm 0.0104$ $0.7885\pm 0.0205$ $0.8782\pm 0.0127$ $0.8804\pm 0.0071$
ContraWR(Yang et al., 2023) 1.6M $0.4384\pm 0.0349$ $0.3912\pm 0.0237$ $0.6893\pm 0.0136$ $0.7851\pm 0.0050$ $0.8719\pm 0.0076$ $0.8743\pm 0.0042$
CNN-Transformer(Peh et al., 2022) 3.2M $0.4087\pm 0.0161$ $0.3815\pm 0.0134$ $0.6854\pm 0.0293$ $0.7898\pm 0.0062$ $0.8781\pm 0.0053$ $0.8781\pm 0.0067$
FFCL(Li et al., 2022) 2.4M $0.3979\pm 0.0104$ $0.3732\pm 0.0188$ $0.6783\pm 0.0120$ $0.7869\pm 0.0054$ $0.8694\pm 0.0091$ $0.8761\pm 0.0044$
ST-Transformer(Song et al., 2021) 3.5M $0.3984\pm 0.0228$ $0.3765\pm 0.0306$ $0.6823\pm 0.0190$ $0.8004\pm 0.0037$ $0.8737\pm 0.0069$ $0.8798\pm 0.0029$
BIOT(Yang et al., 2024) 3.2M $0.4679\pm 0.0354$ $0.4890\pm 0.0407$ $0.7352\pm 0.0236$ $0.7955\pm 0.0047$ $0.8819\pm 0.0046$ $0.8834\pm 0.0041$
LaBraM-Base(Jiang et al., 2024) 5.8M $0.4682\pm 0.0856$ $0.5067\pm 0.0413$ $0.7466\pm 0.0202$ $0.7720\pm 0.0046$ $0.8498\pm 0.0036$ $0.8534\pm 0.0027$
TFM-Token-R 1.8M $\underline{0.4898}\pm 0.0105$ $0.5194\pm 0.0195$ $0.7518\pm 0.0095$ $\underline{0.8033}\pm 0.0021$ $\underline{0.8908}\pm 0.0027$ $\underline{0.8849}\pm 0.0024$
TFM-Token-S 1.9M $0.4708\pm 0.0339$ $\underline{0.5275}\pm 0.0314$ $\underline{0.7538}\pm 0.0152$ $0.7927\pm 0.0044$ $0.8814\pm 0.0095$ $0.8836\pm 0.0052$
TFM-Token 1.9M 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163 0.8152 $\pm$ 0.0014 0.8946 $\pm$ 0.0008 0.8897 $\pm$ 0.0008
1. The best results are bolded, while the second-best are underlined. 2. The number of parameters for LaBraM is only considering their classifier model. The size of their neural tokenizer was 8.6M. 3. Results on CHB-MIT and IIIC Seizure are provided in Table 4 and 5 in Appendix A.1
In this section, we perform comprehensive experiments to evaluate our proposed method to answer the following research questions:
  • Q1: Can a fully discretized EEG tokenization framework outperform baseline methods that rely on continuous embeddings? How crucial is joint frequency-temporal modeling in EEG analysis and tokenization?
  • Q2: How effectively do our learned tokens represent EEG features and capture class-specific characteristics? Do the generated tokens capture relevant frequency information?
  • Q3: Are the tokens generated by our TFM-Tokenizer scalable, and can they enhance the performance of existing EEG foundation models such as LaBraM?
  • Q4: Do tokens produced by TFM-Tokenizer offer interpretability, and do they correspond to distinct, recognizable EEG patterns?
5.1. Experiment Setup
5.1.1. Datasets:
We conducted our evaluation on four EEG datasets:
  • TUH EEG Events (TUEV) (Harati et al., 2015): TUEV is a subset of the TUH EEG Corpus (Obeid and Picone, 2016), which comprises clinical EEG recordings collected at Temple University Hospital between 2002 and 2017. The dataset is annotated for six EEG event types: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
  • TUH Abnormal EEG Corpus (TUAB) (Lopez et al., 2015): TUAB comprises EEG recordings collected at Temple University Hospital, which are labeled for normal and abnormal EEG activity.
  • IIIC Seizure (Jing et al., 2023; Ge et al., 2021): The IIIC Seizure dataset is curated for the detection of six distinct ictal–interictal–injury continuum (IIIC) patterns and is sourced from (Jing et al., 2023; Ge et al., 2021). The annotations include: (1) others (OTH), (2) seizure types (ESZ), (3) lateralized periodic discharge (LPD), (4) generalized periodic discharge (GPD), (5) lateralized rhythmic delta activity (LRDA), and (6) generalized rhythmic delta activity (GRDA).
  • CHB-MIT (Shoeb, 2009): The CHB-MIT dataset is a widely used benchmark for epilepsy seizure detection. It comprises EEG recordings from 23 pediatric subjects with intractable seizures.
5.1.2. Preprocessing:
We follow the preprocessing setup of BIOT (Yang et al., 2024). Unlike LaBraM (Jiang et al., 2024), which utilized 23 channels in the TUEV and TUAB datasets, we adhere to the 16-channel bipolar montage from the international 10–20 system, as used in (Yang et al., 2024). All EEG recordings are resampled to 200 Hz. For TUEV and TUAB, we apply a bandpass filter (0.10.10.10.175757575 Hz) and a notch filter (50 Hz), following the preprocessing pipeline of LaBraM (Jiang et al., 2024). STFT computation of the signals is performed using PyTorch, with detailed parameters provided in Appendix C.2. For training, validation, and test splits, we follow the recommendations from (Yang et al., 2024). Additional details on dataset statistics and splits are provided in Appendix C.1.
5.1.3. Baselines and Metrics:
We evaluated our approach against the baselines from (Yang et al., 2024) as well as the current state-of-the-art methods, including BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024). All baselines were reproduced using their respective open-source GitHub repositories. To ensure a fair comparison, our experiments follow a single-dataset setting for all the baselines. Specifically for BIOT, we conducted their proposed unsupervised pretraining followed by fine-tuning on the same dataset. Similarly, for LaBraM, we used their base model and conducted neural tokenizer training, masked EEG modeling, and fine-tuning within the same dataset. For performance evaluation, we used balanced accuracy, Cohen’s Kappa coefficient, and weighted-F1 score for multi-class classification tasks, while balanced accuracy, AUC-PR, and AUROC were used for binary classification tasks. For TUAB, we used binary cross-entropy loss for fine-tuning, while the cross-entropy loss was applied to the TUEV and IIIC datasets. Given the class imbalance in the CHB-MIT dataset, we employed focal loss for all experiments. All experiments were conducted using five different random seeds, and we report the mean and standard deviation for each metric. Appendix C provides additional details on the experiment settings.
5.2. Q1: Performance and Importance of Joint Frequency-Temporal Modeling
Q1.1 - Performance Evaluation: Table 1 presents EEG event classification results on TUEV and abnormal detection performance on TUAB. Our TFM-Token consistently outperforms all baselines across all metrics. For event-type classification on TUEV, TFM-Token achieves a $5\%$ increase in balanced accuracy over BIOT ($0.4679\rightarrow 0.4943$) and LaBraM ($0.4682\rightarrow 0.4943$). In Cohen’s Kappa, TFM-Token improves by $9\%$ over BIOT ($0.4890\rightarrow 0.5337$) and $5\%$ over LaBraM ($0.5067\rightarrow 0.5337$). In abnormal detection on TUAB, TFM-Token achieves a $5\%$ improvement across all metrics compared to LaBraM. This highlights that our fully discrete tokenization approach surpasses the performance of existing continuous embedding-based approaches. Another advantage of TFM-Token is its reduced model footprint. As shown in Table 1, TFM-Token achieves better results with significantly fewer parameters—a 3-fold reduction compared to LaBraM (5.8M $\rightarrow$ 1.9M) and a 1.5-fold reduction compared to BIOT (3.2M $\rightarrow$ 1.9M). This reduction can be attributed to discrete tokenization approach, which compresses EEG into a token sequence, thereby reducing data complexity.
Q1.2 - Importance of Joint Frequency and Temporal Modeling: To evaluate the importance of joint frequency-temporal modeling, we conducted an ablation study comparing three tokenization variants: (1) TFM-Token-Raw Signal Only (TFM-Token-R), which uses only raw EEG patches $\{x_{i}\}_{i=1}^{N}$ to predict the spectrum $\mathbf{S}$, (2) TFM-Token-STFT Only (TFM-Token-S), and (3) TFM-Token, which jointly models both temporal and frequency features. Masked modeling was applied for token learning in the latter two, with consistent TFM-Encoder training across all variants. As shown in Table 1, all three variants outperform existing baselines. In event classification, TFM-Token-S improves Cohen’s Kappa over TFM-Token-R ($0.5194\rightarrow 0.5275$). However, in abnormal detection, TFM-Token-R achieves a higher AUC-PR ($0.8814\rightarrow 0.8908$). These results indicate that different EEG tasks rely on distinct feature domains, underscoring the necessity of joint modeling. The primary TFM-Token consistently outperforms both single-domain approaches across all settings, further underscoring the importance of joint modeling.
5.3. Q2: EEG Token Quality Analysis and Frequency Learning
Table 2. Token Utilization and class-token uniqueness comparison
Tokenization Method # Params Utilization Class-Token
[…]
TUEV IIIC TUEV IIIC
Neural Tokenizer (LaBraM) 8.6M 21.13 15.25 0.034 0.000
TFM-Tokenizer-R 1.1M 5.29 7.87 0.000 0.000
TFM-Tokenizer-S 1.1M 13.93 11.04 0.004 0.619
TFM-Tokenizer 1.2M 9.78 8.26 2.14 1.429
We study the quality of the EEG tokens learned by our TFM-Tokenizer by analyzing four key aspects: (1) token utilization, (2) class-specific distinctiveness, (3) similar class retrieval, and (4) frequency learning capability. We conducted our analysis using all three TFM-Tokenizer variants and the neural tokenizer from LaBraM (Jiang et al., 2024), testing them on the test splits of both the TUEV and IIIC datasets, which have multiple classes. All tokenizers employed a fixed vocabulary size of $8,192$ tokens for consistency and fair comparison.
[…]
Figure 3a visualizes the class-token uniqueness scores for each class in both datasets. A robust tokenizer should capture class-distinctive tokens across all dataset classes through unsupervised pretraining. To assess this, we computed the geometric mean (GM) of class-token uniqueness scores, as shown in Table 2. Our TFM-Tokenizer reduces token utilization by more than two-fold compared to the neural tokenizer on TUEV ($21.13\%\rightarrow 9.78\%$) and nearly two-fold on IIIC ($15.25\%\rightarrow 8.26\%$). It also significantly improves learning of class-unique tokens compared to neural tokenizer ($0.034\%\rightarrow 2.14\%$on TUEV, $0.0\%\rightarrow 1.429\%$ on IIIC). These results demonstrate that the TFM-Tokenizer captures more compact and useful tokens than the neural tokenizer. Additionally, TFM-Tokenizer achieves a higher class-token uniqueness score across all classes compared to TFM-Tokenizer-R ($0.0\%\rightarrow 1.429\%$ on IIIC) and TFM-Tokenizer-S ($0.619\%\rightarrow 1.429\%$ on IIIC), as depicted in Figure 3a. This further validates joint frequency-temporal modeling in EEG analysis.
Q2.2 - Tokens for Similar-Class Sample Mining: We conducted an EEG signal mining experiment based on similar-class sample retrieval. Given a multi-channel EEG sample, we first obtain its discrete token representation. Using the Jaccard similarity score, we then retrieve the top $K$ most similar samples from the dataset and compute the precision score for correctly retrieving samples of the same class. For this study, we constructed a balanced subset from the IIIC and TUEV datasets and tested all four tokenization methods. The retrieval performance, illustrated in Figure 3b, shows that all TFM-Tokenizer variants significantly outperform neural tokenizer. Notably, TFM-Tokenizer-S and TFM-Tokenizer achieve nearly $60\%$ precision on the TUEV for $K=1$. While the Jaccard similarity measure demonstrates initial feasibility, further research is needed to identify optimal metrics for token-based EEG retrieval.
Figure 3. Analysis of token quality across three TFM-Tokenizer variants and the neural tokenizer. (a) Comparison of class-token uniqueness scores across all classes. (b) Retrieval performance comparison of tokenizers in a similar-class sample mining task.
Q2.3 - Evaluating the Frequency Learning of TFM-Tokenizer Tokens: In this experiment, we compare the frequency and temporal-domain encoders of the TFM-Tokenizer to evaluate their ability to capture diverse frequency features in EEG signals. Specifically, we arrange all tokens in temporal order and perform a discrete Fourier transform on the token sequence. This process decomposes the tokens into frequencies, where each frequency reflects the degree of change between tokens at various scales. Larger changes indicate more diverse token representations. Then, we compute spectral entropy, defined as the normalized Shannon entropy of the amplitude values, to quantify how energy is distributed across the spectrum. Higher spectral entropy means that the model has learned a broader range of frequency features, capturing differences from both large-scale trends and fine details. Figure 4 shows that on the TUEV, TUAB, and CHBMIT datasets, the frequency encoder produces tokens with significantly higher spectral entropy than the temporal encoder. For example, on the TUEV dataset, the frequency encoder achieved an average spectral entropy of 0.26, while the temporal encoder reached only 0.14. This multi-scale sensitivity benefits downstream tasks such as classification, where learning detailed differences in EEG tokens can improve performance.
Figure 4. An analysis of how the proposed frequency and temporal-domain encoders capture frequency features, by using the spectral entropy of the learned token sequences from randomly selected samples. Higher values indicate that the tokens contain richer frequency information.
5.4. Q3: Does TFM-Tokenizer Enhance LaBraM?
Table 3. Performance Comparison of LaBraM with their neural tokenizer vs TFM-Tokenizer
Dataset Tokenizer Performance Metrics
Balanced Acc. Cohen’s Kappa Weighted F1
TUEV Neural Tokenizer $0.4682\pm 0.0856$ $0.5067\pm 0.0413$ $0.7466\pm 0.0202$
TFM-Tokenizer 0.5147±$\pm 0.0174$uparrow 0.5220±$\pm 0.0153$uparrow 0.7533±$\pm 0.0094$uparrow
Balanced Acc. AUC-PR AUROC
TUAB Neural Tokenizer $0.7720\pm 0.0046$ $0.8498\pm 0.0036$ $0.8534\pm 0.0027$
TFM-Tokenizer 0.7765±$\pm 0.0016$uparrow 0.8518±$\pm 0.0051$uparrow 0.8584±$\pm 0.0022$uparrow
To assess the scalability of TFM-Tokenizer, we investigated its ability to enhance an existing EEG foundation model. We selected LaBraM (Jiang et al., 2024), which employs a neural tokenizer solely for pretraining. This setup makes it an ideal candidate for this study. We replaced LaBraM neural tokenizer with TFM-Tokenizer during the masked EEG modeling stage and evaluated its performance on TUEV and TUAB, presented in Table 3. On TUEV, LaBraM with TFM-Tokenizer achieves a $9\%$ increase in balanced accuracy ($0.4682\rightarrow 0.5147$) and a $3\%$ increase in Cohen’s Kappa ($0.5067\rightarrow 0.5220$). On TUAB, TFM-Tokenizer consistently outperforms the neural tokenizer. These results confirm the capability of TFM-Token in enhancing the performance of EEG foundation models. The increase in balanced accuracy suggests that our tokenizer learns more class-discriminative tokens than the neural tokenizer.
Figure 5. Overview of motifs captured by TFM-Tokenizer on the TUEV dataset: (a) presents three samples from the PLED class, while (b) displays three samples from the GPED class.
5.5. Q4: Interpretability of Learned Tokens
We conducted a visual inspection study to determine whether the learned tokens by our TFM-Tokenizer capture and represent distinct time-frequency motifs in EEG signals. Our findings are presented in Figure 5, which illustrates the unique tokens identified by TFM-Tokenizer on the TUEV dataset. Each token represents a spectral window and its corresponding raw EEG patch (1s window with 0.5s overlap). For more precise visualization, we highlight each class’s most frequently occurring tokens, assigning different colors to each token. This analysis confirms that our TFM-Tokenizer effectively captures class-specific distinct EEG patterns, encoding them into discrete tokens. For instance, token $4035$ in the PLED class consistently captures a characteristic drop followed by a rise in EEG signals, maintaining its structure across different samples despite variations in noise, amplitude, and minor shifts within the window. Similarly, tokens such as $5096$ and $3751$ in the GPED class demonstrate the advantages of combining frequency and temporal features, as they remain robust to minor temporal shifts and warping within a window due to emphasizing spectral patterns. However, we also identified certain limitations associated with using fixed windowing and overlapping segments for tokenization. Specifically, when large shifts cause a distinct EEG pattern to be split across two adjacent windows, the model may assign them separate tokens, potentially treating them as distinct patterns.
6. Conclusion
We introduced TFM-Token, a fully discrete tokenization framework consisting of TFM-Tokenizer and TFM-Encoder modules. TFM-Tokenizer effectively encodes time-frequency motifs from single-channel EEG into distinct tokens, which TFM-Encoder utilizes for downstream tasks. Comprehensive evaluations across multiple datasets demonstrate that TFM-Token outperforms state-of-the-art models while using fewer parameters. Our analysis confirms the robustness, class distinctiveness, and interpretability of learned tokens, providing deeper insights into EEG tokenization. These findings lay a strong foundation for future research in developing tokens that extend beyond fixed window segments and enhance capturing temporal-frequency information.
Appendix
[…]
1Contents
Appendix A Additional Results
A.1. Performance on CHB-MIT and IIIC Seizure
Table 4 and 5 presents the performance comparison of TFM-Token with baselines on seizure detection (CHB-MIT) and seizure type classification (IIIC Seizure) tasks. TFM-Token outperforms all baselines across all metrics in both datasets. On the CHB-MIT dataset with a highly imbalanced binary classification task, BIOT is the only baseline with an AUC-PR above $0.25$. However, TFM-Token surpasses BIOT, achieving an $8\%$ improvement in AUC-PR ($0.3127\rightarrow 0.3379$) and a $4.5\%$ increase in AUROC ($0.8456\rightarrow 0.8839$), demonstrating better robustness to class imbalance. For the IIIC Seizure dataset, where the task is to classify 10-second, 16-channel EEG segments into six classes, TFM-Token improves Cohen’s Kappa by $9.5\%$ ($0.4549\rightarrow 0.4985$) and Weighted F1 by $8.5\%$ ($0.5387\rightarrow 0.5847$) over ContraWR, which achieves second best results.
The superior performance of TFM-Token across four EEG datasets shows the promise of a fully discretized framework that has the potential to enhance future EEG foundation models. These results also underscore the importance of capturing both temporal and frequency information, highlighting the critical role of frequency learning in EEG analysis.
Table 4. Seizure detection performance comparison on the CHB-MIT dataset
Models Number CHB-MIT (seizure detection)
of Params Balanced Acc. AUC-PR AUROC
SPaRCNet(Jing et al., 2023) 0.79M $0.5483\pm 0.0164$ $0.1264\pm 0.0112$ $0.7953\pm 0.0402$
ContraWR(Yang et al., 2023) 1.6M $0.6479\pm 0.0423$ $0.2312\pm 0.0316$ $0.8198\pm 0.0197$
CNN-Transformer(Peh et al., 2022) 3.2M $0.5900\pm 0.0429$ $0.2250\pm 0.0512$ $0.8333\pm 0.0326$
FFCL(Li et al., 2022) 2.4M $0.6565\pm 0.0260$ $0.2115\pm 0.0286$ $0.8071\pm 0.0117$
ST-Transformer(Song et al., 2021) 3.5M $0.6717\pm 0.0271$ $0.2133\pm 0.0148$ $0.8657\pm 0.0139$
BIOT(Yang et al., 2024) 3.2M $0.6582\pm 0.0896$ $0.3127\pm 0.0890$ $0.8456\pm 0.0333$
LaBraM-Base(Jiang et al., 2024) 5.8M $0.5035\pm 0.0078$ $0.0959\pm 0.0742$ $0.6624\pm 0.1050$
TFM-Token 1.9M 0.6750 $\pm$ 0.0392 0.3379 $\pm$ 0.0515 0.8839 $\pm$ 0.0173
Table 5. Seizure type classification performance comparison on the IIIC Seizure dataset
Models Number IIIC Seizure (seizure type classification)
of Params Balanced Acc. Cohen’s Kappa Weighted F1
SPaRCNet(Jing et al., 2023) 0.79M $0.5011\pm 0.0286$ $0.4115\pm 0.0297$ $0.4996\pm 0.0262$
ContraWR(Yang et al., 2023) 1.6M $0.5421\pm 0.0123$ $0.4549\pm 0.0166$ $0.5387\pm 0.0138$
CNN-Transformer(Peh et al., 2022) 3.2M $0.5395\pm 0.0144$ $0.4500\pm 0.0165$ $0.5413\pm 0.0176$
FFCL(Li et al., 2022) 2.4M $0.5309\pm 0.0217$ $0.4412\pm 0.0253$ $0.5315\pm 0.0277$
ST-Transformer(Song et al., 2021) 3.5M $0.5093\pm 0.0122$ $0.4217\pm 0.0151$ $0.5217\pm 0.0110$
BIOT(Yang et al., 2024) 3.2M $0.4458\pm 0.0183$ $0.3418\pm 0.0228$ $0.4511\pm 0.0207$
LaBraM-Base(Jiang et al., 2024) 5.8M $0.4736\pm 0.0101$ $0.3716\pm 0.0128$ $0.4765\pm 0.0097$
TFM-Token 1.9M 0.5775 $\pm$ 0.0042 0.4985 $\pm$ 0.0039 0.5847 $\pm$ 0.0050
A.2. Effect of Masked Token Prediction in EEG Tokenization
We conducted an ablation study on TFM-Encoder to assess the impact of masked token prediction pretraining in a fully discretized framework. Using a pretrained TFM-Tokenizer, we compared two approaches: (1) masked token prediction pretraining followed by fine-tuning and (2) direct fine-tuning without pretraining. This experiment was performed on the TUEV dataset across all three TFM-Tokenizer variants, with results summarized in Figure 6. While Cohen’s Kappa and Weighted F1 showed no significant differences between the two approaches, masked token prediction pretraining significantly improved balanced accuracy across all TFM-Tokenizer variants. This suggests that pretraining enhances class-wise prediction consistency by capturing token dependencies and making TFM-Encoder more robust to missing channels or time segments, a common challenge in EEG analysis.
Figure 6. Masked Token Prediction Ablation
A.3. Removing Position Embedding in TFM-Tokenizer Improves Token Learning
Table 6. TFM-Tokenizer Comparison with and w/o Position Embedding (PE) on TUEV Dataset
Method Utilization Uniqueness Balanced Cohen’s Weighted
[…]
TFM-Tokenizer + PE $12.87$ $1.94$ $0.4765\pm 0.038$ $0.5119\pm 0.022$ $0.7457\pm 0.012$
TFM-Tokenizer w/o PE $9.78$ $2.14$ 0.4943 ±$\pm 0.052$ 0.5337 ±$\pm 0.031$ 0.7570 ±$\pm 0.016$
Through our empirical analysis, we found that the performance of our TFM-Token significantly improved when no position embedding was applied to the TFM-Tokenizer. EEG patterns are inherently chaotic and non-stationary, meaning similar motifs can occur at any position within the signal. An ideal tokenizer should be capable of capturing and representing such EEG motifs as distinct tokens without relying on positional information.
We conducted an ablation study comparing the TFM-Tokenizer’s performance with and without position embeddings to critically analyze this phenomenon. The results of this analysis, presented in Table 6, clearly show that the TFM-Tokenizer without position embedding achieves significantly better performance, with an increase of $4\%$ in Cohen’s Kappa ($0.5119\rightarrow 0.5337$).
We further studied the quality of the learned tokens in terms of token utilization and class-uniqueness scores. Token utilization decreased ($12.87\%\rightarrow 9.78\%$) when position embeddings were removed, while the class-token uniqueness score increased ($1.94\%\rightarrow 2.14\%$). This suggests that the TFM-Tokenizer, when using positional encoding, learns different tokens for the same motifs depending on their location in the signal, leading to redundancy. Removing the position embedding allows the TFM-Tokenizer to learn more compact and meaningful tokens without introducing unnecessary data complexities. This improvement is further illustrated in the motifs captured by the TFM-Tokenizer’s tokens in Figure 5 in Section 5.5.
A.4. Ablation on Token Vocabulary Size
To evaluate the impact of token vocabulary size on performance and token learning, we conducted an ablation study by varying the vocabulary size from 256 to 8192 in powers of two. As shown in Figure 7, no monotonic trend was observed for Cohen’s Kappa and Weighted F1 scores. However, balanced accuracy increased with larger vocabulary sizes. Further analysis of token utilization and class-token uniqueness scores is presented in Figure 8. Notably, Figure 8b shows that class-token uniqueness scores increase with vocabulary size, contributing to the improvement in balanced accuracy by enabling learning more unique class-specific tokens.
Figure 7. Token vocabulary size ablation with performance metrics
Figure 8. Token vocabulary size ablation with token utilization and uniqueness
A.5. Ablation on Masking
Table 7. Ablation on masking used for the pretraining of TFM-Tokenizer on TUEV Dataset
Masking Strategy Balanced Acc. Cohen’s Kappa Weighted F1
Random Masking $0.4351\pm 0.0462$ $0.4772\pm 0.0140$ $0.7296\pm 0.0076$
Frequency Bin Masking $0.4673\pm 0.0540$ $0.5193\pm 0.0243$ $0.7536\pm 0.0125$
Frequency Bin $\mathbf{0.4946\pm 0.0392}$ $0.5045\pm 0.0221$ $0.7462\pm 0.0116$
+ Temporal Masking
Frequency Bin $0.4943\pm 0.0516$ $\mathbf{0.5337\pm 0.0306}$ $\mathbf{0.7570\pm 0.0163}$
+ Temporal Masking
+ Symmetric Masking
We conducted an ablation study on masking strategies during TFM-Tokenizer pretraining to assess their impact on performance. Results shown in Table 7 indicate that random masking on the spectrogram $S$ performs poorly compared to other strategies, underscoring the need for effective masking to capture frequency and temporal features from EEG. Frequency bin masking significantly improves performance over random masking, with an $8\%$ increase in Cohen’s Kappa ($0.4772\rightarrow 0.5193$) and a $7\%$ increase in balanced accuracy ($0.4351\rightarrow 0.4673$), highlighting the importance of modeling frequency band dynamics. The addition of temporal masking further boosts balanced accuracy by $5\%$ ($0.4673\rightarrow 0.4946$), underscoring the importance of joint temporal-frequency modeling. However, temporal masking results in a decline in Cohen’s Kappa and Weighted F1, which is then resolved by introducing symmetric masking, achieving the overall best performance.
Appendix B TFM-Token Implementation and Hyperparameter Tuning
Figure 9. TFM-Token Overview
Figure 9 presents an overview of TFM-Token during inference. This section provides additional details on the implementation and training of the framework.
B.1. Training Pipeline:
For all experiments, we follow a single-dataset setting, where all processes in each experiment are conducted within the same dataset. The training process of our framework is as follows: (1) TFM-Tokenizer unsupervised pretraining, (2) TFM-Encoder pretraining using masked token prediction, and finally (3) fine-tuning on the same dataset for downstream tasks.
B.2. Hyperparameter Tuning of TFM-Token
We employed a systematic approach to optimize the hyperparameters of both the TFM-Tokenizer and TFM-Encoder models using Ray Tune [^1] https://docs.ray.io/en/latest/tune/ with the Optuna [^2] https://optuna.org/ search algorithm. Our optimization process followed a three-phase strategy.
In the first phase, we optimized the TFM-Tokenizer architecture by tuning the depth and number of attention heads in the frequency transformer, temporal transformer, and transformer decoder modules to minimize the masked reconstruction loss $\mathcal{L}_{recon}$. This was followed by tuning the training optimizer’s parameters, including learning rate and weight decay. The second phase focused on the TFM-Encoder optimization for the classification task, where we first tuned its architectural parameters (depth and number of heads), followed by training the optimizer’s parameters while keeping the tokenizer frozen. The third phase focused on tuning optimizer parameters for the masked token prediction pretraining of the TFM-Encoder.
To ensure a fair comparison with LaBraM’s neural tokenizer, we maintained a vocabulary size of $8,192$ and an embedding dimension of $64$. For our ablation studies involving raw signal-only and STFT-only variants, we doubled the embedding dimensions of the temporal encoder and frequency patch encoder to match the codebook dimension while maintaining all other parameters same. Detailed hyperparameter configurations for both TFM-Tokenizer and TFM-Encoder are provided in Appendices B.3 and B.4, respectively.
B.3. TFM-Tokenizer Hyperparameters
Table 8. Hyperparameters for TFM-Tokenizer unsupervised pretraining on single-channel setting
Hyperparameter Values
Batch size 256
Optimizer AdamW
Weight decay 0.00001
$\beta_{1}$ 0.9
$\beta_{2}$ 0.99
Learning rate scheduler Cosine
Minimal Learning rate 0.001
Peak Learning rate 0.005
# of Warmup Epochs 10
# of Pretraining Epochs 100
Table 9. Hyperparameters for TFM-Tokenizer
Hyperparameter Values
Temporal Encoder Convolution layer 1 Input Channels 1
Output Dimension 64
Kernel Size 200
Stride 100
Convolution layer 2 Output Dimension 64
Kernel Size 1
Stride 1
Convolution layer 3 Output Dimension 32
Kernel Size 1
Stride 1
Frequency Patch Encoder Convolution layer 1 Input Channels 1
Output Dimension 64
Kernel Size 5
Stride 5
Convolution layer 2 Output Dimension 64
Kernel Size 1
Stride 1
Convolution layer 3 Output Dimension 64
Kernel Size 1
Stride 1
Frequency Transformer Transformer Encoder Layers 2
Embedding Dimension 64
Number of Heads 8
Gated Patchwise Aggregation Output Dimension 32
Kernel Size 5
Stride 5
Temporal Transformer Transformer Encoder Layers 2
Embedding Dimension 64
Number of Heads 8
Token vocabulary (Codebook size) 8192
Transformer Decoder Transformer Encoder Layers 8
Embedding Dimension 64
Number of Heads 8
Linear Decoder 100
B.4. TFM-Encoder Hyperparameters
Table 10. Hyperparameters for TFM-Encoder, its masked token prediction pretraining and downstream finetuning
Hyperparameter Values
Transformer Encoder Layers 4
Embedding Dimension 64
Number of Heads 8
Masked Token Prediction Pretraining
Batch size 512
Optimizer AdamW
Weight decay 0.00001
$\beta_{1}$ 0.9
$\beta_{2}$ 0.99
Learning rate scheduler Cosine
Minimal Learning rate 0.001
Peak Learning rate 0.005
# of Warmup Epochs 5
# of training Epochs 50
Finetuning
Other parameters are the same as above except:
$\beta_{2}$ 0.999
label smoothing (multi-class) 0.1
Appendix C Additional Experiment Details
C.1. Dataset Statistics and Splits
Table 11. Dataset Summary
Dataset # of Recordings # of Samples Duration (s) Task
TUEV $11,914$ $112,491$ $5$ EEG Event Classification
IIIC Seizure $2,689$ $135,096$ $10$ Seizure Type Classification
CHB-MIT $686$ $326,993$ $10$ Seizure Detection
TUAB $2,339$ $409,455$ $10$ Abnoral EEG Detection
This section provides detailed information on the datasets used in our experiments and their respective splits. Table 11 summarizes key statistics, including the number of recordings, the total number of samples after preprocessing, their duration, and the corresponding downstream tasks. For TUEV and TUAB, we utilized the official training and test splits provided by the dataset and further divided the training splits into $80\%$ training and $20\%$ validation sets. We performed a subject-wise split into $60\%$ training, $20\%$ validation, and $20\%$ test on the IIIC Seizure dataset. In the CHB-MIT dataset, we used 1-19 subjects for training, 20-21 for validation, and 22-23 for testing.
C.2. STFT parameters
To extract frequency-domain representations of the EEG, we utilized the STFT function from PyTorch. The recommendations of (Yang et al., 2024) guided our parameter selection and empirical analysis of different configurations to optimize the time-frequency resolution tradeoff. The final parameters are as follows:
Table 12. STFT parameters
Parameter Value Description
FFT size ($n_{\text{fft}},L$) $200$ Number of frequency bins (equal to resampling rate)
Hop length $H$ $100$ Step size for sliding window ($50\%$ overlap)
Window type Hann A smoothing window function to reduce spectral leakage
Output representation Magnitude Only the absolute values of the STFT are retained
Centering False The STFT is computed without implicit zero-padding
One-sided output True Only the positive frequency components are kept

### v2

@@
-Single-Channel EEG Tokenization Through Time-Frequency Modeling
-Abstract.
-We introduce TFM-Tokenizer, a novel tokenization framework tailored for EEG analysis that transforms continuous, noisy brain signals into a sequence of discrete, well-represented tokens for various EEG tasks.
-Conventional approaches typically rely on continuous embeddings and inter-channel dependencies, which are limited in capturing inherent EEG features such as temporally unpredictable patterns and diverse oscillatory waveforms.
-In contrast, we hypothesize that critical time-frequency features can be effectively captured from a single channel.
-By learning tokens that encapsulate these intrinsic patterns within a single channel, our approach yields a scalable tokenizer adaptable across diverse EEG settings.
-We integrate the TFM-Tokenizer with a transformer-based TFM-Encoder, leveraging established pretraining techniques from natural language processing—such as masked token prediction—followed by downstream fine-tuning for various EEG tasks.
-Experiments across four EEG datasets show that TFM-Token outperforms state-of-the-art methods.
-On TUEV, our approach improves balanced accuracy and Cohen’s Kappa by $5\%$ over baselines.
-Comprehensive analysis of the learned tokens demonstrates their ability to capture class-distinctive features, enhance frequency representation, and ability to encode time-frequency motifs into distinct tokens, improving interpretability.
-EEG, Tokenization, Deep learning
-journalyear: 2025copyright: acmcopyrightconference: 2025; ; ccs: Computing methodologies Machine learning approachesccs: Applied computing Health care information systems
-1. Introduction
-Electroencephalograms (EEGs) reveal real-time neuronal activity with millisecond precision, reflecting the responses to various event stimuli.
-This makes EEGs essential not only for fundamental research (Conrad et al., 2019; Li et al., 2021) but also for diverse applications such as sleep staging (Phan et al., 2022; Yang et al., 2023; Pradeepkumar et al., 2024), neurological disease detection (Kotoge et al., 2024; Afzal et al., 2024), and brain-computer interfaces (Precup and Bachman, 2012; Wolf et al., 2022; Li et al., 2023).
-Practical EEG analysis is a complex, multi-step process involving filtering, feature engineering, pattern recognition, annotation, etc.
-Deep learning (DL) models have shown remarkable success in automating EEG analysis across various tasks (Sharma et al., 2022; Tang et al., 2022; Chen et al., 2023).
-These achievements stem from their representation learning capability, projecting noisy EEGs onto a discriminative feature space that captures associations with discrete neurophysiological events.
-Recent neuroscientific literature reports that DL methods successfully match the reliability and accuracy of 20 clinical experts in identifying seizure-related events (Jing et al., 2023).
-Figure 1.
-TFM-Tokenizer tokenizes EEG signals by encoding time-frequency motifs as discrete tokens.
-The neural tokenizer used for LaBraM (Jiang et al., 2024), which is a raw signal-based tokenizer, fails to capture time-frequency variations.
-Our TFM-Tokenizer effectively preserves this information.
-Despite their success, effectively representing EEGs remains a primary focus in the research community.
-Real-world EEG recordings are acquired under diverse scenarios, using different devices, and even for the same task, data formats often vary due to mismatched channels and variable lengths (Yang et al., 2024).
-Unfortunately, most existing methods typically learn representations on a case-by-case basis with specific architectures or fixed channel settings.
-These methods exhibit limited generalization across tasks and poor scalability to different data formats.
-For example, a sleep apnea model is not readily applicable to a sleep staging task, since the former aims to capture anomaly patterns while the latter is to learn discriminative representations (Pradeepkumar et al., 2024).
-Likewise, a sleep staging model trained on central region EEGs cannot directly generalize to frontal EEG data.
-As a result, retraining, elaborate fine-tuning, or model adjustments are often required (Yang et al., 2024), escalating practical costs and further restricting applicability to advanced devices or non-standard data, such as ECoG (Pradeepkumar et al., 2021).
-There is thus an urgent need to develop an EEG analysis method that serves broader research objectives.
-Recently, the transformative impact of large foundation models(Touvron et al., 2023; Achiam et al., 2023) has inspired and elevated EEG representation learning to new heights.
-Several foundation EEG models have been proposed (Duan et al., 2023; Yi et al., 2024; Mohammadi Foumani et al., 2024; Yang et al., 2024), demonstrating both enhanced performance and generalization.
-Researchers often tokenize EEGs into discrete, short-duration snapshots across different data formats and model their dependencies leveraging foundational yet powerful Transformers.
-By leveraging the paradigm of pretraining on large-scale, cross-task datasets, these models can learn universal representation, enabling effective adaptation to downstream tasks without requiring task-specific architectures (Kotoge et al., 2024; Mohammadi Foumani et al., 2024).
-Thanks to unified tokenization and the automated input adaptation of the Transformer, such methods also address challenges such as complex data format reconfigurations (Yang et al., 2024) and labeled data scarcity (Jiang et al., 2024).
-However, this direction remains nascent, and several limitations remain:
-  • Inappropriate Tokenization Representation.
-One reason large language models (LLMs) succeed is their effective tokenization and similar benefits have been demonstrated in image (Van Den Oord et al., 2017) and video (Wang et al., 2024b; Agarwal et al., 2025) tokenization.
-However, existing foundation EEG models generally do not adopt a discrete tokenization paradigm.
-Although some methods claim to provide an EEG “tokenizer,” they typically lack a discrete approach similar to NLP.
-For instance, LaBraM (Jiang et al., 2024) employs a neural tokenizer only during pretraining but relies on raw EEG signals during inference.
-BIOT (Yang et al., 2024) suggests a “biosignal tokenizer” that remains continuous rather than discretized.
-Although time-series tokenization methods have shown promise (Ansari et al., 2024; Rasul et al., 2023), they do not scale well to EEG’s higher sampling rates and other artifacts.
-  • Insufficient Frequency Representation.
-Capturing eventful EEG features, which are characterized by distinct frequencies, is a primary focus of EEG analysis.
-However, tokenizing and modeling raw EEGs often lead to a loss of frequency diversity.
-This frequency representation collapse is a common issue in time-series modeling, as low-frequency components typically dominate the EEG data, biasing models toward lower frequencies while overlooking critical high-frequency features, such as spikes.
-Although LaBraM has recently incorporated frequency reconstruction into its objective function, it still relies on raw EEG signals, inevitably resulting in sub-optimal representation, as shown in Figure 1.
-  • Scalability and Generalization.
-EEG-related tasks vary significantly in channel configurations.
-For example, seizure detection typically utilizes 16 channels, whereas sleep studies often require only 1–2 channels.
-However, existing models are primarily designed for multi-channel settings, heavily relying on cross-channel prediction.
-This design limits their scalability and adaptability to configurations with fewer or even single channels, as well as to varying acquisition setups.
-Therefore, in this paper, we propose TFM-Token, an effective, fully discretized EEG tokenization framework that captures time-frequency motifs from single-channel EEG signals into distinct tokens.
-Technically, our contributions are as follows:
-  • TFM-Tokenizer and TFM-Encoder: We introduce a scalable discrete tokenization framework for EEG, transforming single-channel signals into discrete token sequences akin to NLP models.
-TFM-Tokenizer converts EEG into discrete tokens, and TFM-Encoder utilizes them for downstream tasks.
-  • Joint Modeling of Frequency and Temporal Dynamics: Our tokenizer integrates raw EEG patches with time-frequency representations, using frequency band and temporal masking to capture essential frequency patterns while disentangling temporal variations (Figure 1).
-  • Scable tokenization: Our single-channel approach enables flexible adaptation across EEG tasks and channel configurations.
-TFM-Tokenizer further enhances state-of-the-art EEG models, such as LaBraM (Jiang et al., 2024).
-  • Empirical Validation and Token Quality Analysis: We evaluate our framework on four EEG downstream tasks, demonstrating state-of-the-art performance.
-Beyond performance, we comprehensively analyze token quality, including token visualization, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
-2. Related work
-EEG Representation Learning.
-To learn general representations and address issues of label scarcity in EEG data, self-supervised learning (SSL) has emerged as a prominent paradigm, and existing works can be categorized into two main approaches: contrastive learning and self-prediction.
-Contrastive learning methods, including TS-TCC (Eldele et al., 2021), TF-C (Zhang et al., 2022) and SplitSEE (Kotoge et al., 2024), leverage augmentation or transformation of EEG inputs to learn consistent representations.
-In contrast, self-prediction methods, namely BENDR (Kostas et al., 2021), MAEEG (Chien et al., 2022), BIOT (Yang et al., 2024) and EEG2REP (Mohammadi Foumani et al., 2024), aim to accurately reconstruct masked or corrupted input.
-However, their learning objectives heavily rely on cross-channel prediction to focus on spatial characteristics.
-In contrast, our method emphasizes inherent time-frequency features within a single-channel setting and can adapt to any channel configuration.
-Foundation EEG Models.
-Inspired by the success of foundation models in NLP, recent efforts have sought to develop foundation models for EEG analysis.
-These models can be categorized into decoding and encoder-based methods.
-Decoding-only methods focus on generative tasks like EEG-to-text translation, with representative works including DeWave (Duan et al., 2023), EEG2Text (Liu et al., 2024), and E2T-PTR (Wang et al., 2024a).
-In contrast, encoder-only methods concentrate on fundamental EEG classification tasks and representation learning.
-Notable models include LaBraM (Jiang et al., 2024), BIOT (Yang et al., 2024), BRANT (Zhang et al., 2024), and MMM (Yi et al., 2024).
-Our work aligns with this latter category, focusing on enhancing the representation quality to improve classification performance.
-EEG Tokenization.
-Tokenization has been instrumental in NLP, where discrete subword units have proven to reduce data complexity and improve model performance and interoperability.
-Existing attempts for EEGs include patch-based continuous tokenization, such as BIOT (Yang et al., 2024) and BRANT (Zhang et al., 2024), and vector quantization (VQ)-based methods like DeWave (Duan et al., 2023).
-Patch-based methods do not involve encoding or quantization, leading to unbounded and continuous representations that lack distinctiveness and interpretability.
-In contrast, VQ-based tokenizers, traditionally successful in tokenizing continuous images (Esser et al., 2020), have recently been adapted for EEG by LaBraM (Jiang et al., 2024), However, its tokenizer is utilized only during training and not during inference.
-Conceptually, its primary role is to pre-train classification layers, rather than encoding inputs and reducing data complexity.
-Here, our method is explicitly VQ-based, treating the codebook as a real tokenizer for EEG data.
-Moreover, we enforce each token to capture time-frequency motifs(Xu et al., 2023) in EEG inputs, ensuring a more structured and interpretable representation.
-Frequency Representation Collapse.
-Frequency domain analysis is crucial in EEG and general time series analysis (Elvander and Jakobsson, 2020; Wu et al., 2021, 2023; Woo et al., 2022).
-In real-world signals, time-domain observations inherently mix multiple frequency components, and high-energy, low-frequency signals often dominate the spectrum (Huang Norden E Shen Zheng and H, 1998; Lai et al., 2018).
-As a result, these entangled frequency features makes it difficult for models to distinguish between them (Zhou et al., 2022; Piao et al., 2024).
-Recent studies have shown that these entangled signals can lead to a collapse in the learned frequency representations (Zhi-Qin John Xu et al., 2020; Piao et al., 2024).
-Models tend to overemphasize the dominant low-frequency features while neglecting the high-frequency details.
-This issue can lead to a lack of capturing various EEG waveforms and degenerating data representation (Park and Kim, 2022).
-Motivated by these works, our paper focuses on developing methods to learn diverse, informative frequency features.
-In Section 3, we provide an analysis of our proposed frequency-domain tokenizer and its impact on model performance.
-3. Preliminary
-Figure 2.
-Overview of the TFM-Token framework.
+Tokenizing Single-Channel EEG with Time-Frequency Motif Learning
+Abstract
+Foundation models are reshaping EEG analysis, yet an important problem of EEG tokenization remains a challenge.
+This paper presents TFM-Tokenizer, a novel tokenization framework that learns a vocabulary of time-frequency motifs from single-channel EEG signals and encodes them into discrete tokens.
+We propose a dual-path architecture with time–frequency masking to capture robust motif representations, and it is model-agnostic, supporting both lightweight transformers and existing foundation models for downstream tasks.
+Our study demonstrates three key benefits: Accuracy: Experiments on four diverse EEG benchmarks demonstrate consistent performance gains across both single- and multi-dataset pretraining settings, achieving up to $17\%$ improvement in Cohen’s Kappa over strong baselines.
+Generalization: Moreover, as a plug-and-play component, it consistently boosts the performance of diverse foundation models, including BIOT and LaBraM.
+Scalability: By operating at the single-channel level rather than relying on the strict 10–20 EEG system, our method has the potential to be device-agnostic.
+Experiments on ear-EEG sleep staging, which differs from the pretraining data in signal format, channel configuration, recording device, and task, show that our tokenizer outperforms baselines by $14\%$.
+A comprehensive token analysis reveals strong class-discriminative, frequency-aware, and consistent structure, enabling improved representation quality and interpretability.
+Code is available at https://github.com/Jathurshan0330/TFM-Tokenizer.
+1 Introduction
+Foundation models have revolutionized how machines understand human language, leading to major breakthroughs in natural language processing (NLP) (OpenAI et al., 2024; DeepSeek-AI et al., 2025) and cross-modality tasks such as text-to-image generation (Bordes et al., 2024).
+Inspired by this success, researchers are now advancing a paradigm shift in electroencephalogram (EEG) analysis toward task-agnostic foundation models (Mohammadi Foumani et al., 2024; Yang et al., 2024; Jiang et al., 2024b; Wang et al., 2024a).
+By pretraining on massive, diverse EEG data corpora, these models learn universal representations that generalize well across various downstream tasks.
+Despite substantial recent progress, an important open problem remains: how to design an effective tokenization method for EEG signals.
+Tokenization, a core component in NLP, transforms raw text into meaningful tokens, which reduces data complexity and introduces a helpful inductive bias in foundation models (Gastaldi et al., 2025).
+Typically, tokenization is performed by a learnable function that trains a vocabulary of tokens and statistics from a given corpus.
+However, existing EEG foundation models tokenize signals by directly segmenting continuous EEGs into short-duration tokens, without learning a vocabulary.
+They merely discretize EEG signals, failing to capture statistically grounded representations in a data-driven manner.
+LaBraM (Jiang et al., 2024b) proposes a neural tokenizer to learn data-driven tokens before pretraining.
+However, these tokens primarily serve as training objectives rather than as actual inputs for subsequent model training and are discarded during downstream inference, limiting their reusability.
+As a result, the foundation model is still trained on continuous segment-level embeddings, failing to fully leverage the benefits of tokenization, such as improving the quality of input representations.
+In this paper, we study a novel and critical problem of developing a principled EEG tokenization that seamlessly integrates with various foundation models and enhance downstream performance and generalization.
+Figure 1: (a) Our TFM-Tokenizer converts single-channel EEG into discrete tokens by capturing time-frequency motifs.
+(b) It is adaptable to any different multi-channel settings, (c) can be integrated with existing foundation models to enhance their performance, and (d) enables cross-device scalability.
+Various studies have shown that developing an effective tokenization is a non-trivial task in general, as it is influenced by multiple factors (Schmidt et al., 2024).
+In this paper, we recognize and focus on three key challenges of EEG tokenization.
+1) Tokenization target: real-world EEG recordings exhibit diverse formats due to varying devices, channel configurations, and recording lengths (Yang et al., 2024).
+We argue that tokenizers should be trained and operated at the single-channel level to learn channel-agnostic discrete tokens.
+This design enables flexible adaptation to multi-channel tasks and can generalize to non-standard EEG devices.
+In Section 4.4, we provide scalability experiments on ear-EEG settings.
+2) Token resolution: in NLP, tokenization can be defined at different resolutions (characters, subwords, words), each reflecting different assumptions about semantic granularity.
+However, EEG signals are characterized by diverse oscillatory (e.g., alpha, beta) (Pradeepkumar et al., 2024) and transient patterns (e.g., spikes) (Chen et al., 2022).
+Thus, effective tokens must represent such underlying motifs (Xu et al., 2023) that reflect distinct neural or physiological events.
+However, these motifs are often distorted by noise, amplitude scaling, and temporal warping, making it challenging to design robust EEG tokenization methods.
+3) Tokenization learning objective: EEGs exhibit various temporal variations, manifested as a mixture of low- and high-frequency components that co-occur and are intermixed in complex ways.
+Relying solely on capturing time‑based motifs into discrete tokens risks losing important spectral structure.
+We therefore argue that the tokenization learning objective should incorporate time–frequency representations, enabling tokens to encode more meaningful EEG motifs.
+To tackle these challenges, we propose TFM-Tokenizer, a novel EEG tokenization framework that captures time–frequency motifs from single-channel EEG signals and encodes them into distinct tokens.
+Specifically, 1) Tokenizing EEGs at single-channel: We tokenize single-channel EEG signals into discrete token sequences akin to NLP models, which are then paired with a generic transformer to perform multi-channel modeling using these single-channel tokens.
+Our tokenizer is model-agnostic and can be paired with any downstream model.
+Our experiments confirmed that TFM-Tokenizer can seamlessly integrate with existing foundation models, and further improve their performance (see Figure 1).
+2) Learning motif features as tokens: We introduce a motif learning architecture that encodes time–frequency motifs into tokens through a dual-path encoding design.
+Capturing frequency-band characteristics or compositions is crucial for EEG analysis, and to model such dynamics, we designed a Localized Spectral Window Encoder, which isolates and aggregates information across frequency bands prior to fusion with temporal features.
+3) Explicit time-frequency masking prediction: this learning objective disentangles the entangled time–frequency representations, enabling the model to explicitly learn distinct frequency-specific patterns across time.
+By forcing the model to predict masked regions in both domains, it encourages the tokenizer to discover and encode meaningful neural motifs that are localized in time and frequency.
+Overall, our contributions are summarized as follows:
+  • Formulating Single-Channel EEG Tokenization.
+To our knowledge, we are the first to investigate the problem of learning a discrete token vocabulary that captures time–frequency motifs in single-channel EEG signals from a given corpus and directly utilizes them as inputs for downstream modeling.
+  • Proposing Novel TFM-Token Framework.
+We introduce a single-channel EEG tokenization framework that transforms EEG into a discrete token sequence via TFM-Tokenizer, which is then used by a lightweight transformer model for cross-channel and downstream modeling.
+As shown in Figure 1c, TFM-Tokenizer integrates smoothly with existing models and consistently boosts performance, improving BIOT and LaBraM by approximately $4\%$ on TUEV dataset.
+  • Broad Evaluation across Foundation Models and Devices.
+Extensive experiments across four datasets show that our method outperforms strong baselines, achieving up to a $17\%$ gain over the baseline model on TUEV dataset.
+We also evaluate cross-device scalability on an ear-EEG sleep staging task, using electrodes outside the standard 10–20 EEG system, where our tokenizer outperforms baselines by $14\%$.
+Beyond performance, we comprehensively analyze token quality, including token consistency, class-specific uniqueness, and frequency learning analysis, validating that our learned tokens are informative and interpretable.
+2 Related Work
+EEG Foundation Models and Tokenization Methods.
+Existing EEG foundation models can be categorized into decoding and encoder-based methods.
+Decoding-based methods focus on generative tasks like cross-modal translation  (Duan et al., 2023; Liu et al., 2024; Wang et al., 2024c).
+In contrast, encoder-based methods focus on classification tasks and representation learning.
+Notable models include LaBraM  (Jiang et al., 2024b), BIOT  (Yang et al., 2024), BRANT  (Zhang et al., 2024), and MMM  (Yi et al., 2024).
+Our work aligns with this latter category, aiming to enhance input representations to improve classification performance and generalization across diverse foundation models.
+A parallel question is how to tokenize EEG signals.
+Existing methods primarily adopt segment-based continuous tokenization (Yang et al., 2024; Wang et al., 2024b; Zhang et al., 2024).
+Vector Quantized (VQ) tokenizers (Van Den Oord et al., 2017), which have been successful in tokenizing continuous images (Esser et al., 2020), have recently been adapted for EEG by LaBraM (Jiang et al., 2024b).
+However, in LaBraM, the tokenizer is not designed to represent EEG data and replace raw signals as inputs to foundation models; instead, it mainly serves as a training objective.
+In this paper, we propose a new tokenization framework for EEG signals that encodes inputs into discrete representations and provide a reusable interface for foundation models.
+EEG Motif Learning.
+Motifs are short, recurring patterns with small variability in a time series and may hold predictive or discriminative value (Xu et al., 2023).
+In the EEG domain, motif learning remains largely underexplored, with only a few works such as (Schäfer & Leser, 2022), which focus solely on the temporal domain.
+EEG motifs correspond to neurophysiological events such as oscillatory bursts or transient spikes, which are best characterized by joint temporal-spectral structure.
+Frequency-domain modeling is therefore essential, yet raw time-domain signals often entangle multiple spectral components.
+This can cause models to overemphasize dominant low-frequency rhythms while overlooking informative high-frequency details (Zhi-Qin John Xu et al., 2020; Piao et al., 2024).
+Such bias limits the ability to capture diverse EEG waveforms and degrades representation quality (Park & Kim, 2022).
+To the best of our knowledge, we are the first to propose methods to encode diverse, informative time–frequency motifs as discrete tokens.
+3 Methodology
+3.1 Framework Overview and Forward Process
+Our TFM-Tokenizer framework consists of two major phase, as shown in Figure 2:
+  1. 1.
+    TFM-Tokenizer with Motif Learning.
+The tokenizer is trained in a single-channel, unsupervised setting, capturing key motif features.
+We regard motifs as various waveforms that encode characteristic time–frequency patterns in EEGs.
+To represent these motifs, the tokenizer is composed of four components: (i) a Localized Spectral Window Encoder that extracts frequency patterns within short spectral windows, (ii) a Temporal Encoder that incorporates raw EEG context, (iii) a Temporal Transformer that models dependencies across windows, and (iv) a codebook quantizer that maps embeddings into a discrete vocabulary.
+Therefore, we train a motif-based vocabulary that transforms continuous EEGs into interpretable discrete tokens (Sec.
+3.2).
+  2. 2.
+    Downstream Transformer Model.
+This phase serves as an example to illustrate how a foundation model processes tokenized sequences for downstream tasks such as classification.
+Raw EEGs are first passed through our pretrained tokenizer, where they are converted into discrete tokens that serve as inputs to foundation models.
+Since the tokenizer is model-agnostic, it can be paired with different backbone models.
+In our implementation, we adopt a lightweight Transformer (Vaswani, 2017) with linear attention (Katharopoulos et al., 2020), demonstrating that the tokenizer ($\sim$0.7M parameters) enables strong performance even with a compact model (Sec.
+3.3).
+Overall, we first pretrain the tokenizer to learn a discrete vocabulary of EEG motifs.
+The tokenizer is then frozen, and the downstream Transformer is pretrained with a masked token prediction objective.
+Finally, the downstream Transformer is fine-tuned on target EEG tasks such as classification.
+Figure 2: Overview of our framework.
 (a) TFM-Tokenizer Pretraining: Through dual-path encoding and masked prediction, learns to capture time-frequency motifs into discrete tokens.
-(b) TFM-Encoder Pretraining: Trains on learned EEG tokens using masked token prediction.
-(c) Masking Strategy: A combination of frequency band masking and temporal masking is used for TFM-Tokenizer pretraining.
-(d) Localized Spectral Window Encoder: Processes individual spectral windows from $\mathbf{S}$, extracts frequency band information, and aggregates features across all bands into a single compact embedding per window.
-3.1. Notations and Problem Statements
-EEG Data.
-Let $\mathbf{X}\in\mathbb{R}^{C\times T}$ denote a multi-channel EEG recording with $C$ channels and $T$ time samples.
-For each channel $c\in\{1,...,C\}$, we denote a single channel EEG by $x^{c}\in\mathbb{R}^{T}$.
-To capture complementary structures in both the time and frequency domains, we decompose $x^{c}$ into: (1) time-frequency representation $\mathbf{S}$ and (2) raw EEG patches $\{x_{i}\}_{i=1}^{N}$.
-For simplicity, we omit the channel index; henceforth, $x$ will refer to a single-channel EEG signal unless otherwise specified.
-Short-Time Fourier Transform (STFT).
-To obtain the time-frequency representation, i.e.g, spectrogram, $\mathbf{S}$, we apply a STFT to $x$ using a windowing function $w(.)$ of length $L$ and a hop size $H$:
-(1) $\mathbf{S}(\omega,\tau)=\left|\sum_{l=0}^{L-1}x(\tau H+l)w(l)e^{\frac{-j2\pi\omega l}{L}}\right|$
-where $\omega$ indexes the discrete frequencies and $\tau$ indexes the time segments (i.e., time windows shifted by $H$).
-We retain only the magnitude |.||.|| .
-| to form $\mathbf{S}\in\mathbb{R}^{F\times N}$, where $F$ is the number of frequency bins and $N$ is the number of time windows.
-Problem Statement 1 (EEG Tokenization): Given a single channel EEG $x$, we aim to learn a tokenization function
-$f_{\text{tokenizer}}:\mathbb{R}^{T}\rightarrow\mathcal{V}^{N\times D}$
-where $\mathcal{V}$ is a finite EEG token vocabulary of size $k$, and $D$ is the dimension of each token embedding.
-The tokenizer function $f_{\text{tokenizer}}$ should project $x$ (or transformations) into a sequence of discrete tokens $\{v_{i}\}_{i=1}^{N}$, where each $v_{i}\in\mathcal{V}$.
-These tokens represent various temporal and frequency “motifs”: meaningful EEG patterns characterized by distinct temporal and frequency characteristics.
-Therefore, $\mathcal{V}$ is learnbale from $\mathbf{S}$ and the temporal patches $\{x_{i}\}_{i=1}^{N}$.
-Remark.
-We here hold several expectations for the learned motif tokens.
-First, these tokens are expected to reduce redundancy, noise, and complexity, providing a compact, sparse, and informative representation of EEGs.
-Second, these motifs should effectively capture essential neurophysiological patterns from both temporal and frequency domains.
-Third, the tokens should generalize well across different EEG tasks, enhancing the efficiency and interpretability of the data.
-We set up related research questions to evaluate these expectations in Section.
-5.
-Problem Statement 2 (Multi-Channel EEG Classification): Given EEGs $\mathbf{X}$ and a fixed, learned single-channel tokenizer $f_{\text{tokenizer}}$, we apply $f_{\text{tokenizer}}$ independently to each channel $c$ to obtain a tokenization representation $\Bigl{\{}\{v_{i}^{c}\}_{i=1}^{N}\Bigr{\}}_{c=1}^{C}$.
-Then, these tokens can serve various downstream tasks.
-For classification tasks, they are mapped to labels by:
-$f_{\text{classifier}}:(\mathcal{V}^{D})^{N\times C}\rightarrow\mathbf{Y}$
-where $Y$ is the target labels (e.g., EEG events, seizure types).
-By aggregating and processing the tokens across all channels, $f_{\text{classifier}}$ predicts a label $y\in\mathbf{Y}$.
-Notably, $f_{\text{classifier}}$ can be any downstream-oriented model, and its training is performed separately from the EEG tokenizer $f_{\text{tokenizer}}$.
-4. Methodology
-TFM-Token comprises two components: (1) TFM-Tokenizer: addresses problem statement 1 by converting continuous EEG signals into discrete tokens, capturing key time-frequency features, and (2) TFM-Encoder: tackles problem statement 2 by leveraging these tokens for downstream EEG tasks, such as classification.
-To mitigate the quadratic complexity of standard Transformers (Vaswani, 2017), we employ a linear attention mechanism (Katharopoulos et al., 2020; Wang et al., 2020) across all transformer modules.
-Given input EEGs, TFM-Tokenizer independently tokenizes each channel into discrete token sequences.
-The TFM-Encoder then processes these tokens to perform classification.
-For training TFM-Token, we first conduct an unsupervised pretraining of TFM-Tokenizer in a single-channel setting to encode time-frequency motifs into discrete tokens (see Figure 2a, Sec 4.1).
-The tokenizer is then frozen, and TFM-Encoder undergoes masked token prediction pretraining (see Figure 2b, Sec 4.2), followed by fine-tuning for downstream EEG tasks.
-Time-Frequency discretization.
-The input to TFM-Token consists of discrete time and frequency patches, applied to both raw EEG data and its corresponding spectrogram.
-This approach addresses the transient nature of EEG waveforms, which may decompose the entangled patterns and is foundational for motif learning.
-Specifically, we segment $x$ into time-domain length patches $L$ with the same hop size of $H$.
-This yields a sequence of patches $\{x_{i}\}_{i=1}^{N}$, where $N=\lfloor(T-L)/H\rfloor+1$ is the total number of patches.
-Using the same $L$ and $H$ for both STFT and temporal segmentation, we ensure that each patch $x_{i}$ corresponds one-to-one with a spectral window $\mathbf{S}_{i}$.
-In practice, we obtain the raw EEG patches by configuring the convolutional layers with a kernel size and stride that match the STFT window length $L$ and hop size $H$.
-Additionally, we denote embeddings and masks as $\mathbf{E}$ and $\mathbf{M}$, respectively.
-4.1. Single Channel TFM-Tokenizer
-We introduce the TFM-Tokenizer, a scalable module for tokenizing single-channel EEG signals $x$ by effectively capturing both their temporal and frequency dynamics.
-Our design is inspired by the Vector-Quantized Variational Autoencoder (VQ-VAE)(Van Den Oord et al., 2017), which has been widely adopted for tokenization efforts in other domains such as video processing (Agarwal et al., 2025).
-As illustrated in Figure 2a, the TFM-Tokenizer comprises three primary components: (1) Localized Spectral Window Encoder, (2) Temporal Encoder, and (3) Temporal Transformer.
-At a high level, TFM-Tokenizer adopts a frequency-then-time paradigm.
-First, we isolate each spectral window $\mathbf{S}_{i}$ (for $i=1,…,N$) to capture intra-frequency band dependencies, yielding an aggregated representation for each window (Localized Spectral Window Encoder, see Figure 2d).
-Next, these frequency-based representations are combined with features extracted from the corresponding raw EEG patches $\{x_{i}\}_{i=1}^{N}$ (Temporal Encoder).
-Finally, the combined features are passed through the Temporal Transformer to capture long-range temporal dependencies, and the resulting embeddings are quantized into discrete tokens.
+(b) Masking Strategy: A combination of frequency band masking and temporal masking is used for TFM-Tokenizer pretraining.
+(c) Localized Spectral Window Encoder: Processes individual spectral windows from $\mathbf{S}$, extracts frequency band information, and aggregates features across all bands into a single compact embedding per window.
+(d) Downstream Transformer Encoder Pretraining: Trains on learned EEG tokens using masked token prediction.
+3.2 Single-Channel TFM-Tokenizer with Motif Learning
+TFM-Tokenizer encodes EEGs into discrete motifs tokens through a dual-path frequency–time paradigm (Figure 2a).
+Given a multi-channel EEG $\mathbf{X}\in\mathbb{R}^{C\times T}$, we segment each channel signal $\bm{x}$ into overlapping patches of length $L$ and hop size $H$, yielding $N=\lfloor(T-L)/H\rfloor+1$ patches aligned with spectral windows $\{\mathbf{S}_{i}\}_{i=1}^{N}$.
+To define the pretraining task, masking is applied in both temporal and frequency domains (Figure 2b), where unmasked patches provide context and masked ones are reconstructed.
+Feature learning is performed as follows: each spectral window $\mathbf{S}_{i}$ is encoded by the Localized Spectral Window Encoder (Figure 2c) and fused with raw EEG patch features through a Temporal Encoder.
+A Temporal Transformer then integrates the time–frequency features, and the output embeddings are mapped into a learnable VQ vocabulary, producing motif tokens.
 Localized Spectral Window Encoder.
-Capturing frequency-band characteristics or compositions is crucial for EEG analysis, as the data often contains oscillatory components (e.g., alpha, beta) with varying amplitudes and time.
-Unlike previous studies that project an entire spectral window with a single linear layer (Yang et al., 2024), we propose to patch along the frequency axis, enabling better modeling of cross-frequency dependencies.
-In detail, this encoder consists of three main steps (see Figure 2d):
+Capturing frequency-band characteristics is essential for EEG analysis, as the signals often exhibit oscillatory components (e.g., alpha, beta) with varying amplitudes and temporal dynamics.
+Unlike prior work that projects an entire spectral window through a single linear layer (Yang et al., 2024), we divide the window into patches along the frequency axis, allowing effective modeling of cross-frequency dependencies.
+This process consists of three steps.
   • Frequency Patch Encoder.
 Given a set of spectral windows $\{\mathbf{S}_{i}\}_{i=1}^{N}$, we isolate and divide each spectral window $\mathbf{S}_{i}$ into $P$ non-overlapping patches $\{\mathbf{S}_{(i,p)}\}_{p=1}^{P}$, each spanning $\Delta f$ frequency bins such that $P.\Delta f=F$.
-We then project each frequency patch into a latent space:
-    $e_{(i,p)}=\text{GroupNorm}\left(\text{GeLU}\left(\mathbf{W}_{p}\mathbf{S}_{(i,p)}\right)\right)$
-    where $\mathbf{W}_{p}\in\mathbb{R}^{D\times\Delta f}$ is a learnable matrix.
-  • Frequency Transformer.
-In order to capture the interactions among the frequency patch embeddings $e_{(i,p)}$, we apply a frequency transformer that operates along the frequency axis within each spectral window $\mathbf{S}_{i}$.
-This enables the capturing of intra-frequency band dependencies (e.g., interactions between low and high-frequency bands).
-  • Gated Patchwise Aggregation: In many EEG scenarios, large portions of the frequency spectrum can be irrelevant.
+We then project each frequency patch into a latent space: $e_{(i,p)}=\text{GroupNorm}\left(\text{GeLU}\left(\mathbf{W}_{p}\mathbf{S}_{(i,p)}\right)\right)$ where $\mathbf{W}_{p}\in\mathbb{R}^{D\times\Delta f}$ is the parameter matrix that maps each patch into a $D$-dimensional embedding.
+  • Frequency Transformer. We then apply a frequency transformer that operates along the frequency axis of $\mathbf{S}_{i}$, to model intra-spectral window cross-frequency band dependencies.
+  • Gated Patchwise Aggregation.
+In many EEG scenarios, large portions of the frequency spectrum can be irrelevant.
 For instance, tasks related to sleep primarily focus on frequency bands up to approximately 32 Hz  (Chen et al., 2023).
 Also, the frequencies of interest vary across conditions and tasks.
-To emphasize important frequency patches and suppress the rest, we adopt a gated aggregation mechanism.
-Formally, we first group frequency patch embeddings, then apply gated weighting before concatenating them into a single embedding $\mathbf{E}^{F}_{i}$
-    $\mathbf{E}^{F}_{i}=\text{Concat}\left[\sigma\left(\mathbf{W_{g1}e_{(i,p)}}\right)\mathbf{W_{g2}e_{(i,p)}}\right]$
-    where $\mathbf{W_{g1}},\mathbf{W_{g2}}$ are trainable parameters and $\sigma(\cdot)$ is the element-wise sigmoid function.
-Temporal Encoder.
-To capture temporal dynamics from the raw EEG patches $\{x_{i}\}_{i=1}^{N}$, we perform a linear projection of the patch followed by GELU(Hendrycks and Gimpel, 2016) activation and group normalization, producing temporal embeddings $\{\mathbf{E}^{T}_{i}\}_{i=1}^{N}$.
-Temporal Transformer.
-We form a unified time-frequency embedding by combining each aggregated frequency embedding $\mathbf{E}_{i}^{F}$ with its corresponding temporal embedding $\mathbf{E}_{i}^{T}$ and feed the resulting vectors into a temporal transformer.
-This module attends jointly to time and frequency features across all windows, capturing long-range dependencies.
-The output is then quantized into discrete tokens $\{v_{i}\}_{i=1}^{N}$ using a learnable codebook $\mathcal{V}^{k}$.
-Notably, we do not add any positional encoding as EEG signals are inherently non-stationary and often exhibit chaotic behavior; thus, our goal is to capture their distinctive features without enforcing positional constraints.
-We further provide an ablation study (Appendix A.3) showing the effect of omitting positional encodings.
-Tokenizer Codebook.
-Our tokenizer is designed to capture temporal frequency motifs by applying vector quantization along the time axis, that is, individual patches.
-This is different from conventional quantization methods in computer vision, which typically operate on the embedding dimension of a patch (Van Den Oord et al., 2017; Esser et al., 2020).
-As a result, each token in our vocabulary directly represents a waveform within a specific short duration and can facilitate the retrieval of the timestamp of input EEGs to understand the composition of an EEG.
-This enhances interpretability; for instance, specific codes in the vocabulary are activated in response to the occurrence of particular temporal or frequency patterns.
-To validate this, we conducted a visual inspection study, with results presented in Section 5.5.
-Frequency Making Prediction for Tokenizer Learning.
-The input to the tokenizer is the full frequency band of a temporal patch.
-To facilitate effective frequency learning, we employ a frequency-band and temporal masking strategy during TFM-Tokenizer training.
-We partition $\mathbf{S}$ into $N_{F}=\lfloor\frac{F}{\delta_{f}}\rfloor$ frequency groups of size $\delta_{f}$ and randomly mask a subset based on a predefined mask ratio (e.g., 0.5), forming a frequency-band mask $M_{F}$.
-Simultaneously, a temporal mask $M_{T}$ is applied along the time axis.
-The final mask, $M=M_{F}\land M_{T}$, is used to mask the STFT representation, $S^{M}=M\odot S$, where $\odot$ denotes element-wise multiplication.
-For data augmentation and efficient training, we adopt a symmetric masking approach from (Jiang et al., 2024) by defining $M_{\text{sym}}=1-M$.
-Given $\mathbf{S}^{M}$ and $x$, our TFM-Tokenizer first quantized them into a sequence of distinct tokens, which is then passed through a transformer decoder followed by a linear decoder that predicts the masked regions.
-The tokenizer is trained using a masked reconstruction loss:
-$\mathcal{L}_{\mathrm{rec}}=\sum_{(f,t)\in\mathcal{I}_{\mathrm{mask}}}\!\bigl{\|}\mathbf{S}(f,t)-\hat{\mathbf{S}}(f,t)\bigr{\|}_{2}^{2}$
-where $(f,t)\in\mathcal{I}_{\text{mask}}$ represents masked frequency–time indices and $\hat{\mathbf{S}}$ is the reconstruction of $\mathbf{S}$.
-Alongside the masked reconstruction loss, we incorporate the codebook and commitment losses (Van Den Oord et al., 2017), yielding the total loss:
-$\mathcal{L}_{\mathrm{token}}=\mathcal{L}_{\mathrm{recon}}+\alpha\;\sum_{i}\bigl{\|}\mathrm{sg}[E_{i}]\;-\;v_{i}\bigr{\|}_{2}^{2}+\beta\;\sum_{i}\bigl{\|}E_{i}\;-\;\mathrm{sg}[v_{i}]\bigr{\|}_{2}^{2}$
-where, $\mathrm{sg}[\cdot]$ is the stop-gradient operator, and $\alpha,\beta$ are hyperparameters.
-We also employ exponential moving average strategy(Van Den Oord et al., 2017) for stable codebook updates.
-4.2. Token-Wise TFM-Encoder
-The TFM-Encoder is designed to aggregate tokenized representations across channels and perform downstream classification.
-It consists of two main components: (1) a token-embedding lookup table and (2) linear attention transformer layers.
-Given a multi-channel recording $\mathbf{X}\in\mathbb{R}^{C\times T}$, we apply the fixed (pretrained) TFM-Tokenizer to each channel independently, obtaining discrete token sequences $\Bigl{\{}\{v_{i}^{c}\}_{i=1}^{N}\Bigr{\}}_{c=1}^{C}$.
-Each token is mapped to its embedding via a lookup table initialized with the TFM-Tokenizer ’s learned codebook weights.
-We flatten the embeddings across channels and incorporate channel and position embeddings to encode respective information.
-We then prepend a [CLS] token (Devlin, 2018), the output of which is designed to capture the overall representation of the sequence for downstream tasks.
-The resulting embeddings are then processed by linear attention transformers in TFM-Encoder.
-Token Prediction for TFM-Encoder Learning.
-We adopt a strategy akin to masked language modeling.
-We randomly mask tokens across multiple channels and time steps and train the model to predict these masked tokens via a cross-entropy loss.
+To emphasize important frequency patches and suppress the rest, we adopt a gated aggregation mechanism to obtain a embedding for each $S_{i}$: $\mathbf{E}^{F}_{i}=\text{Concat}\left[\sigma\left(\mathbf{W_{g1}e_{(i,p)}}\right)\mathbf{W_{g2}e_{(i,p)}}\right]$ where $\mathbf{W_{g1}},\mathbf{W_{g2}}$ are trainable parameters and $\sigma(\cdot)$ is the element-wise sigmoid function.
+Temporal Encoder and Temporal Transformer.
+To capture temporal dynamics from raw EEG patches $\{x_{i}\}_{i=1}^{N}$, each patch is projected linearly, followed by GELU activation and group normalization, producing temporal embeddings $\{\mathbf{E}^{T}_{i}\}_{i=1}^{N}$.
+Each aggregated frequency embedding $\mathbf{E}_{i}^{F}$ is then concatenated with its corresponding temporal embedding $\mathbf{E}_{i}^{T}$, and the resulting sequence is processed by a temporal Transformer.
+This module integrates time and frequency features across NN EEG patches, enabling the modeling of long-range dependencies.
+Finally, the outputs $\mathbf{Z}_{i}$ are quantized into discrete tokens using a learnable vocabulary $\mathcal{V}^{k}$.
+Notably, we omit positional encoding because EEG signals are inherently non-stationary and often exhibit chaotic dynamics; our objective is to capture distinctive features without enforcing positional constraints (see Appendix C.6).
+VQ Tokenizer Vocabulary.
+Our vocabulary is based on the discrete codebook of Vector-Quantized Variational Autoencoders (VQ-VAE).
+We perform vector quantization to fused embedding $\mathbf{Z}_{i}$ that enables the vocabulary to capture time–frequency motifs as discrete tokens, supporting timestamp-level retrieval and improving EEG interpretability.
+Formally, given $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{N}$, each $\mathbf{z}_{i}$ is mapped to the closest code in the codebook $\mathcal{V}=\{\mathbf{v}_{1},\dots,\mathbf{v}_{K}\}$ by nearest-neighbor search.
+$q(\mathbf{z}_{i})=\arg\min_{\mathbf{v}_{k}\in\mathcal{V}}\|\mathbf{z}_{i}-\mathbf{v}_{k}\|_{2}^{2}.$
+where $K$ denotes the number of latent vectors in the codebook and defines a $K$-way discrete categorical distribution.
+Each patch $z_{i}$ is mapped to its nearest code entry $v_{i}$.
+As a result, given a single-channel EEG $\mathbf{X}^{c}$, TFM-Tokenizer generates a sequence of NN tokens $\{v_{i}\}_{i=1}^{N}$.
+Frequency Masking Prediction for Tokenizer Learning
+We employ a joint frequency–temporal masking strategy for TFM-Tokenizer training.
+The spectrogram $\mathbf{S}$ is partitioned along the frequency axis into $N_{F}=\lfloor F/\delta_{f}\rfloor$ groups of size $\delta_{f}$, and random frequency-band masks $M_{F}$ and temporal masks $M_{T}$ are applied to obtain the masked input $\mathbf{S}^{M}$.
+Following  (Jiang et al., 2024b), we further adopt symmetric masking for data augmentation and training stability.
+The overall objective combines masked reconstruction and vocabulary loss:
+$\mathcal{L}_{\mathrm{token}}=\sum_{(f,t)}\!\bigl\|\mathbf{S}(f,t)-\hat{\mathbf{S}}(f,t)\bigr\|_{2}^{2}+\alpha\;\sum_{i}\bigl\|\mathrm{sg}[E_{i}]\;-\;v_{i}\bigr\|_{2}^{2}+\beta\;\sum_{i}\bigl\|E_{i}\;-\;\mathrm{sg}[v_{i}]\bigr\|_{2}^{2}$
+where $\hat{\mathbf{S}}$ is the reconstruction, $\mathrm{sg}[\cdot]$ is the stop-gradient operator, and $\alpha,\beta$ are hyperparameters.
+We also apply exponential moving average updates for stable codebook training.
+3.3 Downstream Transformer Training
+We employ a lightweight transformer model to aggregate tokenized representations across channels, learn cross-channel dependencies and perform downstream tasks.
+It consists of a token-embedding lookup table (initialized from the VQ codebook) followed by linear attention transformer layers.
+Given a multi-channel recording $\mathbf{X}\in\mathbb{R}^{C\times T}$, the pretrained TFM-Tokenizer produces token sequences $\Bigl\{\{v_{i}^{c}\}_{i=1}^{N}\Bigr\}_{c=1}^{C}$ for each channel $c$ independently.
+We flatten the token embeddings across channels and incorporate channel and position embeddings.
+An addtional class token is prepended (Devlin, 2018), and the sequence is processed by transformer layers.
+In order to pretrain the model and enable the model to learn intra and cross-channel dependencies of tokens, we adopt a strategy akin to masked language modeling.
+We first randomly mask tokens across multiple channels and time steps and then train the model to predict these masked tokens via a cross-entropy loss.
 Along with representation learning, this approach enhances robustness to missing or corrupted data, common in real-world EEG systems where channels or time segments may be dropped or noisy.
-Finally, the TFM-Encoder is finetuned to downstream tasks.
-5. Experiments and Results
-Table 1. EEG classification performance comparison on TUEV and TUAB datasets.
-Models Number TUEV (event type classification) TUAB (abnormal detection)
-of Params Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC
-SPaRCNet(Jing et al., 2023) 0.79M $0.4161\pm 0.0262$ $0.4233\pm 0.0181$ $0.7024\pm 0.0104$ $0.7885\pm 0.0205$ $0.8782\pm 0.0127$ $0.8804\pm 0.0071$
-ContraWR(Yang et al., 2023) 1.6M $0.4384\pm 0.0349$ $0.3912\pm 0.0237$ $0.6893\pm 0.0136$ $0.7851\pm 0.0050$ $0.8719\pm 0.0076$ $0.8743\pm 0.0042$
-CNN-Transformer(Peh et al., 2022) 3.2M $0.4087\pm 0.0161$ $0.3815\pm 0.0134$ $0.6854\pm 0.0293$ $0.7898\pm 0.0062$ $0.8781\pm 0.0053$ $0.8781\pm 0.0067$
-FFCL(Li et al., 2022) 2.4M $0.3979\pm 0.0104$ $0.3732\pm 0.0188$ $0.6783\pm 0.0120$ $0.7869\pm 0.0054$ $0.8694\pm 0.0091$ $0.8761\pm 0.0044$
-ST-Transformer(Song et al., 2021) 3.5M $0.3984\pm 0.0228$ $0.3765\pm 0.0306$ $0.6823\pm 0.0190$ $0.8004\pm 0.0037$ $0.8737\pm 0.0069$ $0.8798\pm 0.0029$
-BIOT(Yang et al., 2024) 3.2M $0.4679\pm 0.0354$ $0.4890\pm 0.0407$ $0.7352\pm 0.0236$ $0.7955\pm 0.0047$ $0.8819\pm 0.0046$ $0.8834\pm 0.0041$
-LaBraM-Base(Jiang et al., 2024) 5.8M $0.4682\pm 0.0856$ $0.5067\pm 0.0413$ $0.7466\pm 0.0202$ $0.7720\pm 0.0046$ $0.8498\pm 0.0036$ $0.8534\pm 0.0027$
-TFM-Token-R 1.8M $\underline{0.4898}\pm 0.0105$ $0.5194\pm 0.0195$ $0.7518\pm 0.0095$ $\underline{0.8033}\pm 0.0021$ $\underline{0.8908}\pm 0.0027$ $\underline{0.8849}\pm 0.0024$
-TFM-Token-S 1.9M $0.4708\pm 0.0339$ $\underline{0.5275}\pm 0.0314$ $\underline{0.7538}\pm 0.0152$ $0.7927\pm 0.0044$ $0.8814\pm 0.0095$ $0.8836\pm 0.0052$
-TFM-Token 1.9M 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163 0.8152 $\pm$ 0.0014 0.8946 $\pm$ 0.0008 0.8897 $\pm$ 0.0008
+Finally, the transformer model is finetuned for downstream tasks.
+4 Experiments and Results
+4.1 Experiment Setup
+Datasets: We evaluated our method on four EEG datasets.
+(1) TUEV (Harati et al., 2015): A subset of the TUH EEG Corpus (Obeid & Picone, 2016), containing clinical EEG recordings annotated for six event types: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
+(2) TUAB (Lopez et al., 2015): Also from Temple University Hospital, labeled for normal and abnormal EEG activity.
+(3) CHB-MIT (Shoeb, 2009): A widely used benchmark for epilepsy seizure detection, comprising EEG recordings from 23 pediatric subjects with intractable seizures.
+(4) IIIC Seizure (Jing et al., 2023; Ge et al., 2021): Designed for detecting six ictal–interictal–injury continuum (IIIC) patterns, including others (OTH), electrographic seizures (ESZ), lateralized periodic discharges (LPD), generalized periodic discharges (GPD), lateralized rhythmic delta activity (LRDA), and generalized rhythmic delta activity (GRDA).
+- Scalability Validation.
+In this paper, we provided a scalability experiment to evalute the usability of our tokenizer across different EEG devices.
+Since our tokenizer is trained in a single-channel setting, it can naturally be applied to recordings from non-standard devices.
+Therefore, we evaluated on the Ear-EEG Sleep Monitoring (EESM23) (Bjarke Mikkelsen et al., 2025; Tabar et al., 2024) dataset, which contains ear-EEG sleep recordings from 10 subjects.
+Detailed dataset statistics, splits, and preprocessing procedures are provided in Appendix B.1B.2, and  B.3.
+Baselines: We evaluated our approach against the baselines from  Yang et al.
+(2024) and recent state-of-the-art methods, including BIOT, LaBraM, NeuroLM, and EEGPT.
+We adopted the best results reported in BIOT, except for the IIIC Seizure dataset, where we re-evaluated the methods due to a sample size mismatch.
+Experiments were conducted under two settings: (1) Single-dataset setting: pretraining and finetuning on the same single dataset, and (2) Multiple dataset setting: pretraining on four EEG datasets.
+For BIOT, we reproduced their unsupervised pretraining and finetuning pipeline in the single-dataset setting (denoted BIOT) to enable a fair comparison, as their vanilla BIOT variant does not include pretraining.
+Similarly, we reproduced LaBraM by training its neural tokenizer, performing masked EEG modeling, and finetuning within the same dataset (LaBraM).
+Since our focus is on EEG tokenization rather than full foundation modeling, we reproduced LaBraM under the multiple dataset setting using the previously mentioned four EEG datasets (denoted LaBraM).
+This was necessary to ensure a fair comparison because the original LaBraM used a substantially larger pretraining corpus.
+Additional experiment details are provided in Appendix B.4 and  B.5.
+4.2 How Does TFM-Tokenizer Compare to Existing Baselines?
+Table 1 reports results on TUEV (event classification) and TUAB (abnormal detection), while Table 2 summarizes performance on IIIC-Seizure (seizure type classification) and CHB-MIT (seizure detection).
+Our TFM-Tokenizer paired with a downstream transformer consistently outperforms all baselines in both experiment settings.
+On the challenging six-class event-type classification task in TUEV, it achieves a $5\%$ gain in Cohen’s Kappa in the single-dataset setting and a notable $17\%$ improvement ($0.5273\rightarrow 0.6189$) in the multi-dataset setting over the next best baseline.
+On IIIC-Seizure, which is another six-class classification task, TFM-Tokenizer improves Cohen’s Kappa by $36\%$ over the next best baseline LaBraM ($0.3658\rightarrow 0.4979$, p = 1.5e-4) in multiple dataset settings, demonstrating the strong capability of our tokenizer in modeling class-discriminative features for complex clinical EEG tasks.
+Additionally, it is worth noting that TFM-Tokenizer achieves better performance with fewer parameters, being 3 times smaller than LaBraM and 1.5 times smaller than BIOT.
+The ability to achieve best performance with low model size can be attributed to our tokenization approach, which compresses the EEG into a token sequence, thereby reducing data complexity.
+Notably, the TFM-Tokenizer is paired with a lightweight transformer comprising only $\sim$0.7M parameters.
+Table 1: Performance comparison on TUEV and TUAB datasets.
+[…]
+Table 2: Performance comparison on IIIC Seizure and CHB-MIT datasets.
+[…]
 1.
-The best results are bolded, while the second-best are underlined.
+The best and second-best results for each dataset setting are bolded and underlined, respectively.
 2.
@@
 3.
-Results on CHB-MIT and IIIC Seizure are provided in Table 4 and 5 in Appendix A.1
-In this section, we perform comprehensive experiments to evaluate our proposed method to answer the following research questions:
-  • Q1: Can a fully discretized EEG tokenization framework outperform baseline methods that rely on continuous embeddings?
-How crucial is joint frequency-temporal modeling in EEG analysis and tokenization?
-  • Q2: How effectively do our learned tokens represent EEG features and capture class-specific characteristics? Do the generated tokens capture relevant frequency information?
-  • Q3: Are the tokens generated by our TFM-Tokenizer scalable, and can they enhance the performance of existing EEG foundation models such as LaBraM?
-  • Q4: Do tokens produced by TFM-Tokenizer offer interpretability, and do they correspond to distinct, recognizable EEG patterns?
-5.1. Experiment Setup
-5.1.1. Datasets:
-We conducted our evaluation on four EEG datasets:
-  • TUH EEG Events (TUEV) (Harati et al., 2015): TUEV is a subset of the TUH EEG Corpus (Obeid and Picone, 2016), which comprises clinical EEG recordings collected at Temple University Hospital between 2002 and 2017.
-The dataset is annotated for six EEG event types: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
-  • TUH Abnormal EEG Corpus (TUAB) (Lopez et al., 2015): TUAB comprises EEG recordings collected at Temple University Hospital, which are labeled for normal and abnormal EEG activity.
-  • IIIC Seizure (Jing et al., 2023; Ge et al., 2021): The IIIC Seizure dataset is curated for the detection of six distinct ictal–interictal–injury continuum (IIIC) patterns and is sourced from (Jing et al., 2023; Ge et al., 2021).
-The annotations include: (1) others (OTH), (2) seizure types (ESZ), (3) lateralized periodic discharge (LPD), (4) generalized periodic discharge (GPD), (5) lateralized rhythmic delta activity (LRDA), and (6) generalized rhythmic delta activity (GRDA).
-  • CHB-MIT (Shoeb, 2009): The CHB-MIT dataset is a widely used benchmark for epilepsy seizure detection. It comprises EEG recordings from 23 pediatric subjects with intractable seizures.
-5.1.2. Preprocessing:
+$\star$ indicates reproduced in a single dataset setting and $\dagger$ indicates pretraining on 4 EEG datasets.
+4.3 Can TFM-Tokenizer Improve Existing Foundation Models?
+To evaluate the generalizability of TFM-Tokenizer, we integrated it into two representative EEG foundation models, BIOT and LaBraM, under both single- and multi-dataset settings.
+For BIOT, we replaced raw EEG inputs with token embeddings while following the original training protocol.
+For LaBraM, we substituted its neural tokenizer with ours during masked EEG modeling.
+As shown in Figure 3, our method consistently improves performance on TUEV, IIIC, and CHB-MIT, achieving gains of at least $3\%$ in most cases.
+LaBraM notably underperforms on CHB-MIT in the single-dataset setting, yet integrating our tokenizer yields a $147\%$ improvement in AUC-PR, demonstrating its effectiveness in capturing class-discriminative features in data-scarce scenarios.
+These results highlight the broad applicability of TFM-Tokenizer across architectures and its capacity to enhance diverse EEG foundation models.
+Figure 3: Performance comparison of existing foundation models with and without integration of TFM-Tokenizer on the TUEV, IIIC, and CHB-MIT datasets.
+For each dataset, the first three bars show single-dataset pretraining and the latter three show multi-dataset pretraining.
+Percentage values above each bar indicate the relative performance gain achieved by incorporating TFM-Tokenizer.
+4.4 Does TFM-Tokenizer Scale to Other Brain-signal Types / Devices?
+Table 3: Scalability experiments results on EESM23.
+Models Ear-EEG (Sleep Staging)
+Balanced Acc. Cohen’s Kappa Weighted F1
+BIOT 0.3858 $\pm$ 0.0085 0.3406 $\pm$ 0.0096 0.4888 $\pm$ 0.0124
+[…]
+TFM-Tokenizer $\dagger$ 0.4148 $\pm$ 0.0209 0.3883 $\pm$ 0.0233 0.5174 $\pm$ 0.0141
+In order to assess the scalability of TFM-Tokenizer beyond the modalities and tasks seen during pretraining, we evaluate its performance on the EESM23 ear-EEG dataset (Bjarke Mikkelsen et al., 2025) for sleep staging, a task, brain signal modality, acquisition system, number of channels and channel configuration entirely distinct from those in the pretraining set.
+Specifically, we only finetune pretrained models (our method, BIOT, and LaBraM) on the EESM23 dataset using only $\sim$8K labeled training samples.
+EEGPT was not scalable in this setting due to its reliance on a fixed EEG channel layout for spatial embeddings (Wang et al., 2024a).
+As shown in Table 3, TFM-Tokenizer demonstrates strong generalization, outperforming both baselines (p = 0.02) in this out-of-domain setting.
+4.5 How Important are Frequency and Temporal Modeling for EEG Tokenization?
+To evaluate the importance of joint frequency–temporal modeling, we conducted an ablation study with three tokenization variants: (1) TFM-Tokenizer-R, which uses only raw EEG patches to predict the masked spectrogram; (2) TFM-Tokenizer-S, which uses only the spectrogram as input; and (3) TFM-Tokenizer, which jointly models both domains.
+Masked modeling was applied for token learning in the latter two.
+On TUEV (Figure 4a), TFM-Tokenizer-S achieves higher Cohen’s Kappa than TFM-Tokenizer-R, while TFM-Tokenizer-R yields better AUC-PR in abnormal detection (Appendix Figure 6).
+These results show that different EEG tasks rely on different feature domains, underscoring the need for joint modeling, where TFM-Tokenizer consistently outperforms both variants.
+4.6 How Effective are TFM-Tokenizer tokens?
+We evaluate the quality of EEG tokens learned by our tokenizer across four aspects: (1) class-specific distinctiveness, (2) token consistency, (3) frequency learning capability, and (4) token utilization (results in Appendix C.1).
+For this analysis, we compare all three TFM-Tokenizer variants with the neural tokenizer from LaBraM, using the test splits of TUEV and IIIC, which both contain multiple classes.
+To ensure fairness, all tokenizers employ a fixed vocabulary size of $8192$.
+Results on TUEV are shown in Figure 4b–c, with additional results for other datasets provided in the Appendix.
+Class-Token uniqueness.
+To assess whether tokenizers capture class-specific motifs, we define the Class-Token Uniqueness Score as # $\frac{\#\text{ Unique Tokens in Class}}{\#\text{ Tokens Utilized by Class}}\times 100\%.$ This metric quantifies how well a tokenizer assigns distinctive tokens to each class.
+Figure 4b shows the scores for TUEV, where a robust tokenizer should yield high distinctiveness across all classes through unsupervised pretraining.
+TFM-Tokenizer consistently achieves higher scores than its variants and LaBraM’s neural tokenizer, indicating that it produces more compact and informative token representations and validating the benefit of joint frequency–temporal modeling in EEG analysis.
+Class-wise Token Consistency Analysis.
+We conduct a retrieval-based EEG signal mining experiment to evaluate token consistency within the same class, using similar-class sample retrieval (see Figure 4c).
+Given a multi-channel EEG sample, we first obtain its discrete token representation.
+Using the Jaccard similarity score, we then retrieve the top $K$ most similar samples from the dataset and compute the precision score for correctly retrieving samples of the same class.
+For this study, we constructed a balanced subset from the IIIC and TUEV datasets and tested all four tokenization methods.
+Results show that all TFM-Tokenizer variants significantly outperform the neural tokenizer.
+Among all variants, our method yields the best retrieval performance, reflecting better token consistency.
+Notably, TFM-Tokenizer-S and TFM-Tokenizer achieve nearly $60\%$ precision on the TUEV for $K=1$.
+While the Jaccard similarity measure demonstrates initial feasibility, further work is needed to identify optimal metrics.
+Nonetheless, the results suggest that EEG tokens can support the identification of similar pairs, with potential applications in contrastive learning.
+Figure 4: (a) Frequency and temporal token encoder ablation on TUEV. (b) Comparison of class-token uniqueness scores across all classes and (c) Class-wise token consistency analysis.
+4.7 Do the Learned Tokens Capture Meaningful EEG Motifs?
+Figure 5: Overview of motifs captured by TFM-Tokenizer on TUEV: (a) three samples from the PLED class and (b) three samples from the GPED.
+We perform a small-scale qualitative analysis to examine whether TFM-Tokenizer captures meaningful time–frequency motifs in EEG signals.
+Figure 5 shows some representative tokens learned by our method on the TUEV dataset.
+Each token represents a spectral window and its corresponding raw EEG patch (1s window with 0.5s overlap).
+For clarity, we highlight the most frequent tokens per class using distinct colors.
+Periodic Lateralized Epileptiform Discharges (PLEDs) are periodic patterns consisting of sharp waves or spikes followed by a slow wave, occurring every 1-2s (Pohlmann-Eden et al., 1996).
+Token 4035 consistently captures this characteristic waveform across different samples in the PLED class, despite variations in noise, amplitude, and minor temporal shifts.
+This confirms that our TFM-Tokenizer can capture class-specific physiologically meaningful EEG motifs into discrete tokens.
+Similarly, tokens such as $5096$ and $3751$ in the GPED class highlight the benefit of joint time–frequency modeling, as they remain robust to minor temporal shifts and warping within a window due to emphasizing spectral patterns.
+However, we found limitations associated with using fixed windowing for tokenization, as large patterns or shifts may cause splits across windows, leading to separate token assignments and misinterpretation as distinct events.
+5 Conclusion
+In this paper, we presented TFM-Tokenizer, a model-agnostic tokenization framework that encodes single-channel EEG into discrete tokens by capturing time–frequency motifs.
+Our study demonstrated three key benefits: (i) Accuracy: By accurately extracting single-channel features, our tokenizer enabled stronger representations and surpassed competitive baselines across four EEG benchmarks.
+(ii) Generalization: As a plug-and-play component, our method consistently boosted the performance of existing foundation models, showing its broad applicability.
+(iii) Scalability: Because it operates at the single-channel level rather than depending on the strict 10–20 EEG system, our method readily extended to ear-EEG sleep staging tasks, validating its cross-device scalability.
+Furthermore, analyses confirmed the class distinctiveness, consistency, and interpretability of the learned tokens, providing deeper insights into EEG tokenization.
+We hope this work will inspire the development of more robust tokenization frameworks and advance scalable, generalizable EEG foundation models across diverse modalities, devices, and tasks.
+6 Reproducibility statement
+To support the reproducibility of our work, we provide our complete source code and pretrained model weights at https://github.com/Jathurshan0330/TFM-Tokenizer.
+The repository includes scripts for data preprocessing, loading, and model training to reproduce our results presented in this paper.
+In the main text, Section 4.1 outlines our experimental setup, including descriptions of the dataset and baselines.
+Additional implementation details, such as dataset statistics, preprocessing steps, ear-EEG-specific processing, evaluation metrics, and baseline configurations, are provided in Appendix B.1B.2B.3B.4, and B.5.
+The Appendix also includes extended experiments across multiple datasets, including frequency learning analysis (Appendix C.1), cross-dataset generalization studies (Appendix C.3), additional results on improving foundation models (Appendix C.4), and further ablation studies.
+We have made every effort to ensure that our work can be easily reproduced by the community.
+Appendix
+Contents
+Appendix A Problem Formulation
+EEG Data.
+Let $\mathbf{X}\in\mathbb{R}^{C\times T}$ denote a multi-channel EEG recording with $C$ channels and $T$ time samples.
+Each channel $x^{c}\in\mathbb{R}^{T}$ is decomposed into (1) raw patches $\{x_{i}\}_{i=1}^{N}$ and (2) corresponding time-frequency representation windows $\{\mathbf{S}_{i}\}_{i=1}^{N}$, where NN is the number of time windows.
+For simplicity, we omit the channel index and refer to $x$ as a single-channel EEG signal unless stated otherwise.
+To obtain the time-frequency representation, i.e., spectrogram, $\mathbf{S}$, we apply the short-time Fourier transform (STFT) to $x$ using a windowing function $w(.)$ of length $L$ and a hop size $H$.
+Short-Time Fourier Transform (STFT).
+To obtain the time-frequency representation, i.e.g, spectrogram, $\mathbf{S}$, we apply a STFT to $x$ using a windowing function $w(.)$ of length $L$ and a hop size $H$:
+$\mathbf{S}(\omega,\tau)=\left|\sum_{l=0}^{L-1}x(\tau H+l)w(l)e^{\frac{-j2\pi\omega l}{L}}\right|$ (1)
+where $\omega$ indexes the discrete frequencies and $\tau$ indexes the time segments (i.e., time windows shifted by $H$).
+We retain only the magnitude $|.|$ to form $\mathbf{S}\in\mathbb{R}^{F\times N}$, where $F$ is the number of frequency bins and NN is the number of time windows.
+Problem Statement 1 (EEG Tokenization): Given a single channel EEG $x$, we aim to learn a tokenization function $f_{\textbf{tokenizer}}:\mathbb{R}^{T}\rightarrow\mathcal{V}^{N\times D}$, that maps $x$ (or transformations) to a sequence of discrete tokens $\{v_{i}\}_{i=1}^{N}$, where each token is from a learnable EEG token vocabulary $\mathcal{V}$ of size $k$ and embedding size of $D$.
+These tokens should represent various time-frequency “motifs” derived from both $x_{i}$} and $\mathbf{S}_{i}$}.
+Therefore, $\mathcal{V}$ is learnable from $\mathbf{S}$ and the temporal patches $\{x_{i}\}_{i=1}^{N}$.
+Remark.
+We here hold several expectations for the learned motif tokens.
+First, these tokens are expected to reduce redundancy, noise, and complexity, providing a compact, sparse, and informative representation of EEGs.
+Second, these motifs should capture key neurophysiological patterns from both temporal and frequency domains.
+Third, the tokens should generalize well across different EEG tasks.
+[…]
+Appendix B Additional Experiment Details
+B.1 Dataset Statistics and Splits
+Table 4: Evaluation Dataset Summary
+Dataset # of Recordings # of Samples Duration (s) Task
+TUEV $11,914$ $112,491$ 55 EEG Event Classification
+IIIC Seizure $2,689$ $135,096$ $10$ Seizure Type Classification
+CHB-MIT $686$ $326,993$ $10$ Seizure Detection
+TUAB $2,339$ $409,455$ $10$ Abnoral EEG Detection
+EESM23 $120$ $14,509$ $30$ Ear-EEG based Sleep Staging
+This section provides detailed information on the datasets used in our experiments and their respective splits.
+Table 4 summarizes key statistics, including the number of recordings, the total number of samples after preprocessing, their duration, and the corresponding downstream tasks.
+For TUEV and TUAB, we utilized the official training and test splits provided by the dataset and further divided the training splits into $80\%$ training and $20\%$ validation sets.
+We performed a subject-wise split into $60\%$ training, $20\%$ validation, and $20\%$ test on the IIIC Seizure dataset.
+In the CHB-MIT dataset, we used 1-19 subjects for training, 20-21 for validation, and 22-23 for testing.
+For the out-of-distribution evaluation on the ear-EEG EESM23 (Bjarke Mikkelsen et al., 2025) dataset, we followed a subject-wise split, where subjects 1–6 were used for fine-tuning, 7–8 for validation, and 9–10 for testing.
+B.2 Preprocessing
 We follow the preprocessing setup of BIOT  (Yang et al., 2024).
-Unlike LaBraM (Jiang et al., 2024), which utilized 23 channels in the TUEV and TUAB datasets, we adhere to the 16-channel bipolar montage from the international 10–20 system, as used in (Yang et al., 2024).
+We adhere to the 16-channel bipolar montage from the international 10–20 system, as used in  (Yang et al., 2024).
 All EEG recordings are resampled to 200 Hz.
-For TUEV and TUAB, we apply a bandpass filter (0.10.10.10.175757575 Hz) and a notch filter (50 Hz), following the preprocessing pipeline of LaBraM (Jiang et al., 2024).
-STFT computation of the signals is performed using PyTorch, with detailed parameters provided in Appendix C.2.
+For TUEV and TUAB, we apply a bandpass filter (0.10.17575 Hz) and a notch filter (50 Hz), following the preprocessing pipeline of LaBraM  (Jiang et al., 2024b).
+We then segment the recordings according to the provided annotations and preprocessing guidelines.
+STFT computation of the signals is performed using PyTorch, with detailed parameters provided in Appendix B.6.
 For training, validation, and test splits, we follow the recommendations from  (Yang et al., 2024).
-Additional details on dataset statistics and splits are provided in Appendix C.1.
-5.1.3. Baselines and Metrics:
-We evaluated our approach against the baselines from (Yang et al., 2024) as well as the current state-of-the-art methods, including BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024).
-All baselines were reproduced using their respective open-source GitHub repositories.
-To ensure a fair comparison, our experiments follow a single-dataset setting for all the baselines.
-Specifically for BIOT, we conducted their proposed unsupervised pretraining followed by fine-tuning on the same dataset.
-Similarly, for LaBraM, we used their base model and conducted neural tokenizer training, masked EEG modeling, and fine-tuning within the same dataset.
-For performance evaluation, we used balanced accuracy, Cohen’s Kappa coefficient, and weighted-F1 score for multi-class classification tasks, while balanced accuracy, AUC-PR, and AUROC were used for binary classification tasks.
-For TUAB, we used binary cross-entropy loss for fine-tuning, while the cross-entropy loss was applied to the TUEV and IIIC datasets.
-Given the class imbalance in the CHB-MIT dataset, we employed focal loss for all experiments.
-All experiments were conducted using five different random seeds, and we report the mean and standard deviation for each metric.
-Appendix C provides additional details on the experiment settings.
-5.2. Q1: Performance and Importance of Joint Frequency-Temporal Modeling
-Q1.1 - Performance Evaluation: Table 1 presents EEG event classification results on TUEV and abnormal detection performance on TUAB.
-Our TFM-Token consistently outperforms all baselines across all metrics.
-For event-type classification on TUEV, TFM-Token achieves a $5\%$ increase in balanced accuracy over BIOT ($0.4679\rightarrow 0.4943$) and LaBraM ($0.4682\rightarrow 0.4943$).
-In Cohen’s Kappa, TFM-Token improves by $9\%$ over BIOT ($0.4890\rightarrow 0.5337$) and $5\%$ over LaBraM ($0.5067\rightarrow 0.5337$).
-In abnormal detection on TUAB, TFM-Token achieves a $5\%$ improvement across all metrics compared to LaBraM.
-This highlights that our fully discrete tokenization approach surpasses the performance of existing continuous embedding-based approaches.
-Another advantage of TFM-Token is its reduced model footprint.
-As shown in Table 1, TFM-Token achieves better results with significantly fewer parameters—a 3-fold reduction compared to LaBraM (5.8M $\rightarrow$ 1.9M) and a 1.5-fold reduction compared to BIOT (3.2M $\rightarrow$ 1.9M).
-This reduction can be attributed to discrete tokenization approach, which compresses EEG into a token sequence, thereby reducing data complexity.
-Q1.2 - Importance of Joint Frequency and Temporal Modeling: To evaluate the importance of joint frequency-temporal modeling, we conducted an ablation study comparing three tokenization variants: (1) TFM-Token-Raw Signal Only (TFM-Token-R), which uses only raw EEG patches $\{x_{i}\}_{i=1}^{N}$ to predict the spectrum $\mathbf{S}$, (2) TFM-Token-STFT Only (TFM-Token-S), and (3) TFM-Token, which jointly models both temporal and frequency features.
-Masked modeling was applied for token learning in the latter two, with consistent TFM-Encoder training across all variants.
-As shown in Table 1, all three variants outperform existing baselines.
-In event classification, TFM-Token-S improves Cohen’s Kappa over TFM-Token-R ($0.5194\rightarrow 0.5275$).
-However, in abnormal detection, TFM-Token-R achieves a higher AUC-PR ($0.8814\rightarrow 0.8908$).
-These results indicate that different EEG tasks rely on distinct feature domains, underscoring the necessity of joint modeling.
-The primary TFM-Token consistently outperforms both single-domain approaches across all settings, further underscoring the importance of joint modeling.
-5.3. Q2: EEG Token Quality Analysis and Frequency Learning
-Table 2. Token Utilization and class-token uniqueness comparison
+We adopt a window length of 11s with $0.5$s overlap to segment EEG signals during training and inference, following prior work for consistency (Yang et al., 2024).
+B.3 Ear-EEG Preprocessing
+We follow the preprocessing guidelines of Tabar et al.
+(2021) for the EESM-23 ear-EEG dataset, which includes four channels (RB, RT, LB, LT).
+A bandpass filter (0.10.1100100 Hz) and a 50Hz notch filter are applied.
+Each patients perform certain tasks before sleep.
+To isolate sleep segments, we crop each session from the onset of annotated sleep scoring, segment the signal into 30-second epochs, and discard corrupted segments.
+B.4 Evaluation Metrics
+For evaluation, we used balanced accuracy, Cohen’s kappa coefficient, and weighted F1 for multi-class classification, and balanced accuracy, AUROC, and AUC-PR for binary classification.
+During finetuning, we employed binary cross-entropy loss for TUAB, cross-entropy loss for TUEV and IIIC, and focal loss for CHB-MIT due to class imbalance.
+All experiments were conducted using five different random seeds, and we report the mean and standard deviation.
+B.5 Additional details on baselines
+All baselines were reproduced using their official open-source repositories.
+LaBraM’s primary contribution lies in large-scale EEG pretraining using over 2,500 hours of data (Jiang et al., 2024b), whereas our focus is on developing an effective EEG tokenizer.
+To ensure a fair comparison, we reproduced LaBraM using its official repository under our dataset and experimental settings.
+For EEGPT, we report the published results for the 4.7M model on TUEV and TUAB (Wang et al., 2024a).
+Since results on CHB-MIT and IIIC-Seizure were not available, we used the official pretrained weights and fine-tuned the model on these tasks.
+B.6 STFT parameters
+Table 5: STFT parameters
+Parameter Value Description
+FFT size ($n_{\text{fft}},L$) $200$ Number of frequency bins (equal to resampling rate)
+Hop length $H$ $100$ Step size for sliding window ($50\%$ overlap)
+Window type Hann A smoothing window function to reduce spectral leakage
+Output representation Magnitude Only the absolute values of the STFT are retained
+Centering False The STFT is computed without implicit zero-padding
+One-sided output True Only the positive frequency components are kept
+To extract frequency-domain representations of the EEG, we utilized the STFT function from PyTorch.
+The recommendations of Yang et al.
+(2024) guided our parameter selection and empirical analysis of different configurations to optimize the trade-off between time-frequency resolution.
+The final parameters are as follows:
+Appendix C Extended Experiment Results
+C.1 Additional Results on Token Quality Analysis and Frequency Learning
+Figure 6: (a) Frequency and temporal token encoder ablation on TUAB.
+(b) & (c) presents Analysis of token quality across three TFM-Tokenizer variants and the neural tokenizer on IIIC.
+(b) Comparison of class-token uniqueness scores across all classes and (c) Class-wise token consistency analysis
+Table 6: Token Utilization and class-token uniqueness comparison
 Tokenization Method # Params Utilization Class-Token
-[…]
+$\%$ Uniqueness (GM) $\%$
 TUEV IIIC TUEV IIIC
@@
 TFM-Tokenizer 1.2M 9.78 8.26 2.14 1.429
-We study the quality of the EEG tokens learned by our TFM-Tokenizer by analyzing four key aspects: (1) token utilization, (2) class-specific distinctiveness, (3) similar class retrieval, and (4) frequency learning capability.
-We conducted our analysis using all three TFM-Tokenizer variants and the neural tokenizer from LaBraM (Jiang et al., 2024), testing them on the test splits of both the TUEV and IIIC datasets, which have multiple classes.
-All tokenizers employed a fixed vocabulary size of $8,192$ tokens for consistency and fair comparison.
-[…]
-[…]
-Figure 3a visualizes the class-token uniqueness scores for each class in both datasets.
-A robust tokenizer should capture class-distinctive tokens across all dataset classes through unsupervised pretraining.
-To assess this, we computed the geometric mean (GM) of class-token uniqueness scores, as shown in Table 2.
+Figure 7: An analysis of how the proposed frequency and temporal-domain encoders capture frequency features, by using the spectral entropy of the learned token sequences from randomly selected samples.
+Higher values indicate that the tokens contain richer frequency information.
+In this section, we present more results on token quality analysis, specifically focusing on token utilization and frequency learning capability of our tokenizer.
+Additional token uniqueness and consistency experiments on IIIC dataset is presented in Figure 6b and c.
+Token utilization: Token utilization ($\%$) score was calculated as the percentage of unique tokens activated from the total available vocabulary size.
+Additionally, we computed the geometric mean (GM) of class-token uniqueness scores along with the utilization score, and the results are presented in Table 6.
 Our TFM-Tokenizer reduces token utilization by more than two-fold compared to the neural tokenizer on TUEV ($21.13\%\rightarrow 9.78\%$) and nearly two-fold on IIIC ($15.25\%\rightarrow 8.26\%$).
-It also significantly improves learning of class-unique tokens compared to neural tokenizer ($0.034\%\rightarrow 2.14\%$on TUEV, $0.0\%\rightarrow 1.429\%$ on IIIC).
-These results demonstrate that the TFM-Tokenizer captures more compact and useful tokens than the neural tokenizer.
-Additionally, TFM-Tokenizer achieves a higher class-token uniqueness score across all classes compared to TFM-Tokenizer-R ($0.0\%\rightarrow 1.429\%$ on IIIC) and TFM-Tokenizer-S ($0.619\%\rightarrow 1.429\%$ on IIIC), as depicted in Figure 3a.
-This further validates joint frequency-temporal modeling in EEG analysis.
-Q2.2 - Tokens for Similar-Class Sample Mining: We conducted an EEG signal mining experiment based on similar-class sample retrieval.
-Given a multi-channel EEG sample, we first obtain its discrete token representation.
-Using the Jaccard similarity score, we then retrieve the top $K$ most similar samples from the dataset and compute the precision score for correctly retrieving samples of the same class.
-For this study, we constructed a balanced subset from the IIIC and TUEV datasets and tested all four tokenization methods.
-The retrieval performance, illustrated in Figure 3b, shows that all TFM-Tokenizer variants significantly outperform neural tokenizer.
-Notably, TFM-Tokenizer-S and TFM-Tokenizer achieve nearly $60\%$ precision on the TUEV for $K=1$.
-While the Jaccard similarity measure demonstrates initial feasibility, further research is needed to identify optimal metrics for token-based EEG retrieval.
-Figure 3.
-Analysis of token quality across three TFM-Tokenizer variants and the neural tokenizer.
-(a) Comparison of class-token uniqueness scores across all classes.
-(b) Retrieval performance comparison of tokenizers in a similar-class sample mining task.
-Q2.3 - Evaluating the Frequency Learning of TFM-Tokenizer Tokens: In this experiment, we compare the frequency and temporal-domain encoders of the TFM-Tokenizer to evaluate their ability to capture diverse frequency features in EEG signals.
+It also significantly improves learning of class-unique tokens compared to the neural tokenizer ($0.034\%\rightarrow 2.14\%$ on TUEV, $0.0\%\rightarrow 1.429\%$ on IIIC).
+Evaluating the Frequency Learning of TFM-Tokenizer Tokens: In this experiment, we compare the frequency and temporal-domain encoders of the TFM-Tokenizer to evaluate their ability to capture diverse frequency features in EEG signals.
 Specifically, we arrange all tokens in temporal order and perform a discrete Fourier transform on the token sequence.
@@
 This multi-scale sensitivity benefits downstream tasks such as classification, where learning detailed differences in EEG tokens can improve performance.
-Figure 4.
-An analysis of how the proposed frequency and temporal-domain encoders capture frequency features, by using the spectral entropy of the learned token sequences from randomly selected samples.
-Higher values indicate that the tokens contain richer frequency information.
-5.4. Q3: Does TFM-Tokenizer Enhance LaBraM?
-Table 3. Performance Comparison of LaBraM with their neural tokenizer vs TFM-Tokenizer
-Dataset Tokenizer Performance Metrics
-Balanced Acc. Cohen’s Kappa Weighted F1
-TUEV Neural Tokenizer $0.4682\pm 0.0856$ $0.5067\pm 0.0413$ $0.7466\pm 0.0202$
-TFM-Tokenizer 0.5147±$\pm 0.0174$uparrow 0.5220±$\pm 0.0153$uparrow 0.7533±$\pm 0.0094$uparrow
+C.2 Additional results on Frequency and Temporal Modeling for EEG Tokenization
+Table 7: Ablation study on input representation to TFM-Tokenizer
+Models TUEV (event type classification) TUAB (abnormal detection)
+Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC
+TFM-Tokenizer-R $\underline{0.4898}\pm 0.0105$ $0.5194\pm 0.0195$ $0.7518\pm 0.0095$ $\underline{0.8033}\pm 0.0021$ $\underline{0.8908}\pm 0.0027$ $\underline{0.8849}\pm 0.0024$
+TFM-Tokenizer-S $0.4708\pm 0.0339$ $\underline{0.5275}\pm 0.0314$ $\underline{0.7538}\pm 0.0152$ $0.7927\pm 0.0044$ $0.8814\pm 0.0095$ $0.8836\pm 0.0052$
+TFM-Tokenizer 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163 0.8152 $\pm$ 0.0014 0.8946 $\pm$ 0.0008 0.8897 $\pm$ 0.0008
+1. The best results are bolded, while the second-best are underlined.
+In Table 7 we provide detailed results of our ablation study discussed under Section 4.5.
+C.3 Token Generalization Assessment through Cross-Dataset Experiments
+Table 8: Cross dataset generalizability experiments under single dataset settings
+Testing Tokenizer MTP Performance Metrics
+Dataset Dataset Dataset Balanced Acc. Cohen’s Kappa Weighted F1
+TUEV TUEV TUEV 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163
+IIIC TUEV 0.4722 $\pm$ 0.0578 0.4990 $\pm$ 0.0237 0.7380 $\pm$ 0.0137
+IIIC 0.4291 $\pm$ 0.0235 0.5195 $\pm$ 0.0200 0.7534 $\pm$ 0.0100
+TUAB TUEV 0.4651 $\pm$ 0.0449 0.5925 $\pm$ 0.0249 0.7847 $\pm$ 0.0136
+TUAB 0.5252 $\pm$ 0.0431 0.6187 $\pm$ 0.0285 0.8018 $\pm$ 0.0138
+CHB-MIT TUEV 0.4979 $\pm$ 0.0444 0.5995 $\pm$ 0.0225 0.7885 $\pm$ 0.0122
+CHB-MIT 0.5898 $\pm$ 0.0192 0.6591 $\pm$ 0.0106 0.8196 $\pm$ 0.0045
+To evaluate the robustness of our tokenizer, we conducted cross-dataset experiments under two settings: (1) fixing the tokenizer and performing masked token prediction (MTP) & finetuning on a different target dataset and (2) fixing the tokenizer and MTP, followed by finetuning TFM-Encoder only on the target dataset.
+Results are presented in Table 8, which demonstrates strong generalizability, with our TFM-Tokenizer achieving the best performance on TUEV when pretrained on CHBMIT—outperforming the best-reported result in four dataset settings.
+These findings highlight the potential of our tokenizer as a foundation for a scalable, universal EEG tokenizer.
+C.4 Additional Results on TFM-Tokenizer Improving Existing Foundation Models
+Table 9: Performance comparison of LaBraM and BIOT with and w/o our TFM-Tokenizer.
+Dataset Exp. Method Performance Metrics
+Setting Balanced Acc. Cohen’s Kappa Weighted F1
+TUEV Single BIOT 0.4679 $\pm$ 0.0354 0.4890 $\pm$ 0.0407 0.7352 $\pm$ 0.0236
+BIOT-TFM 0.4228 $\pm$ 0.0162 0.5230 $\pm$ 0.0226 $\uparrow$ 0.7490 $\pm$ 0.0114 $\uparrow$
+LaBraM 0.4682 $\pm$ 0.0856 0.5067 $\pm$ 0.0413 0.7466 $\pm$ 0.0202
+LaBraM-TFM 0.5147 $\pm$ 0.0174 $\uparrow$ 0.5220 $\pm$ 0.0153 $\uparrow$ 0.7533 $\pm$ 0.0094 $\uparrow$
+Multiple BIOT 0.5281 $\pm$ 0.0225 0.5273 $\pm$ 0.0249 0.7492 $\pm$ 0.0082
+BIOT-TFM 0.5530 $\pm$ 0.0089 $\uparrow$ 0.5464 $\pm$ 0.0137 $\uparrow$ 0.7625 $\pm$ 0.0069 $\uparrow$
+LaBraM 0.5550 $\pm$ 0.0403 0.5175 $\pm$ 0.0339 0.7450 $\pm$ 0.0194
+LaBraM-TFM 0.5541 $\pm$ 0.0316 0.5367 $\pm$ 0.0281 $\uparrow$ 0.7567 $\pm$ 0.0165 $\uparrow$
+IIIC Single BIOT 0.4458 $\pm$ 0.0183 0.3418 $\pm$ 0.0228 0.4511 $\pm$ 0.0207
+BIOT-TFM 0.4633 $\pm$ 0.0078 $\uparrow$ 0.3663 $\pm$ 0.0103 $\uparrow$ 0.4689 $\pm$ 0.0090 $\uparrow$
+LaBraM 0.4736 $\pm$ 0.0101 0.3716 $\pm$ 0.0128 0.4765 $\pm$ 0.0097
+LaBraM-TFM 0.4814 $\pm$ 0.0075 $\uparrow$ 0.3795 $\pm$ 0.0091 $\uparrow$ 0.4841 $\pm$ 0.0062 $\uparrow$
+Multiple BIOT 0.4414 $\pm$ 0.0035 0.3362 $\pm$ 0.0040 0.4483 $\pm$ 0.0033
+BIOT-TFM 0.5050 $\pm$ 0.0037 $\uparrow$ 0.4098 $\pm$ 0.0052 $\uparrow$ 0.5139 $\pm$ 0.0025 $\uparrow$
+LaBraM 0.4736 $\pm$ 0.0037 0.3658 $\pm$ 0.0033 0.4708 $\pm$ 0.0015
+LaBraM-TFM 0.4782 $\pm$ 0.0065 $\uparrow$ 0.3737 $\pm$ 0.0076 $\uparrow$ 0.4790 $\pm$ 0.0082 $\uparrow$
 Balanced Acc. AUC-PR AUROC
-TUAB Neural Tokenizer $0.7720\pm 0.0046$ $0.8498\pm 0.0036$ $0.8534\pm 0.0027$
-TFM-Tokenizer 0.7765±$\pm 0.0016$uparrow 0.8518±$\pm 0.0051$uparrow 0.8584±$\pm 0.0022$uparrow
-To assess the scalability of TFM-Tokenizer, we investigated its ability to enhance an existing EEG foundation model.
-We selected LaBraM (Jiang et al., 2024), which employs a neural tokenizer solely for pretraining.
-This setup makes it an ideal candidate for this study.
-We replaced LaBraM neural tokenizer with TFM-Tokenizer during the masked EEG modeling stage and evaluated its performance on TUEV and TUAB, presented in Table 3.
-On TUEV, LaBraM with TFM-Tokenizer achieves a $9\%$ increase in balanced accuracy ($0.4682\rightarrow 0.5147$) and a $3\%$ increase in Cohen’s Kappa ($0.5067\rightarrow 0.5220$).
-On TUAB, TFM-Tokenizer consistently outperforms the neural tokenizer.
-These results confirm the capability of TFM-Token in enhancing the performance of EEG foundation models.
-The increase in balanced accuracy suggests that our tokenizer learns more class-discriminative tokens than the neural tokenizer.
-Figure 5. Overview of motifs captured by TFM-Tokenizer on the TUEV dataset: (a) presents three samples from the PLED class, while (b) displays three samples from the GPED class.
-5.5. Q4: Interpretability of Learned Tokens
-We conducted a visual inspection study to determine whether the learned tokens by our TFM-Tokenizer capture and represent distinct time-frequency motifs in EEG signals.
-Our findings are presented in Figure 5, which illustrates the unique tokens identified by TFM-Tokenizer on the TUEV dataset.
-Each token represents a spectral window and its corresponding raw EEG patch (1s window with 0.5s overlap).
-For more precise visualization, we highlight each class’s most frequently occurring tokens, assigning different colors to each token.
-This analysis confirms that our TFM-Tokenizer effectively captures class-specific distinct EEG patterns, encoding them into discrete tokens.
-For instance, token $4035$ in the PLED class consistently captures a characteristic drop followed by a rise in EEG signals, maintaining its structure across different samples despite variations in noise, amplitude, and minor shifts within the window.
-Similarly, tokens such as $5096$ and $3751$ in the GPED class demonstrate the advantages of combining frequency and temporal features, as they remain robust to minor temporal shifts and warping within a window due to emphasizing spectral patterns.
-However, we also identified certain limitations associated with using fixed windowing and overlapping segments for tokenization.
-Specifically, when large shifts cause a distinct EEG pattern to be split across two adjacent windows, the model may assign them separate tokens, potentially treating them as distinct patterns.
-6. Conclusion
-We introduced TFM-Token, a fully discrete tokenization framework consisting of TFM-Tokenizer and TFM-Encoder modules.
-TFM-Tokenizer effectively encodes time-frequency motifs from single-channel EEG into distinct tokens, which TFM-Encoder utilizes for downstream tasks.
-Comprehensive evaluations across multiple datasets demonstrate that TFM-Token outperforms state-of-the-art models while using fewer parameters.
-Our analysis confirms the robustness, class distinctiveness, and interpretability of learned tokens, providing deeper insights into EEG tokenization.
-These findings lay a strong foundation for future research in developing tokens that extend beyond fixed window segments and enhance capturing temporal-frequency information.
-Appendix
-[…]
-1Contents
-Appendix A Additional Results
-A.1. Performance on CHB-MIT and IIIC Seizure
-Table 4 and 5 presents the performance comparison of TFM-Token with baselines on seizure detection (CHB-MIT) and seizure type classification (IIIC Seizure) tasks.
-TFM-Token outperforms all baselines across all metrics in both datasets.
-On the CHB-MIT dataset with a highly imbalanced binary classification task, BIOT is the only baseline with an AUC-PR above $0.25$.
-However, TFM-Token surpasses BIOT, achieving an $8\%$ improvement in AUC-PR ($0.3127\rightarrow 0.3379$) and a $4.5\%$ increase in AUROC ($0.8456\rightarrow 0.8839$), demonstrating better robustness to class imbalance.
-For the IIIC Seizure dataset, where the task is to classify 10-second, 16-channel EEG segments into six classes, TFM-Token improves Cohen’s Kappa by $9.5\%$ ($0.4549\rightarrow 0.4985$) and Weighted F1 by $8.5\%$ ($0.5387\rightarrow 0.5847$) over ContraWR, which achieves second best results.
-The superior performance of TFM-Token across four EEG datasets shows the promise of a fully discretized framework that has the potential to enhance future EEG foundation models.
-These results also underscore the importance of capturing both temporal and frequency information, highlighting the critical role of frequency learning in EEG analysis.
-Table 4. Seizure detection performance comparison on the CHB-MIT dataset
-Models Number CHB-MIT (seizure detection)
-of Params Balanced Acc. AUC-PR AUROC
-SPaRCNet(Jing et al., 2023) 0.79M $0.5483\pm 0.0164$ $0.1264\pm 0.0112$ $0.7953\pm 0.0402$
-ContraWR(Yang et al., 2023) 1.6M $0.6479\pm 0.0423$ $0.2312\pm 0.0316$ $0.8198\pm 0.0197$
-CNN-Transformer(Peh et al., 2022) 3.2M $0.5900\pm 0.0429$ $0.2250\pm 0.0512$ $0.8333\pm 0.0326$
-FFCL(Li et al., 2022) 2.4M $0.6565\pm 0.0260$ $0.2115\pm 0.0286$ $0.8071\pm 0.0117$
-ST-Transformer(Song et al., 2021) 3.5M $0.6717\pm 0.0271$ $0.2133\pm 0.0148$ $0.8657\pm 0.0139$
-BIOT(Yang et al., 2024) 3.2M $0.6582\pm 0.0896$ $0.3127\pm 0.0890$ $0.8456\pm 0.0333$
-LaBraM-Base(Jiang et al., 2024) 5.8M $0.5035\pm 0.0078$ $0.0959\pm 0.0742$ $0.6624\pm 0.1050$
-TFM-Token 1.9M 0.6750 $\pm$ 0.0392 0.3379 $\pm$ 0.0515 0.8839 $\pm$ 0.0173
-Table 5. Seizure type classification performance comparison on the IIIC Seizure dataset
-Models Number IIIC Seizure (seizure type classification)
-of Params Balanced Acc. Cohen’s Kappa Weighted F1
-SPaRCNet(Jing et al., 2023) 0.79M $0.5011\pm 0.0286$ $0.4115\pm 0.0297$ $0.4996\pm 0.0262$
-ContraWR(Yang et al., 2023) 1.6M $0.5421\pm 0.0123$ $0.4549\pm 0.0166$ $0.5387\pm 0.0138$
-CNN-Transformer(Peh et al., 2022) 3.2M $0.5395\pm 0.0144$ $0.4500\pm 0.0165$ $0.5413\pm 0.0176$
-FFCL(Li et al., 2022) 2.4M $0.5309\pm 0.0217$ $0.4412\pm 0.0253$ $0.5315\pm 0.0277$
-ST-Transformer(Song et al., 2021) 3.5M $0.5093\pm 0.0122$ $0.4217\pm 0.0151$ $0.5217\pm 0.0110$
-BIOT(Yang et al., 2024) 3.2M $0.4458\pm 0.0183$ $0.3418\pm 0.0228$ $0.4511\pm 0.0207$
-LaBraM-Base(Jiang et al., 2024) 5.8M $0.4736\pm 0.0101$ $0.3716\pm 0.0128$ $0.4765\pm 0.0097$
-TFM-Token 1.9M 0.5775 $\pm$ 0.0042 0.4985 $\pm$ 0.0039 0.5847 $\pm$ 0.0050
-A.2. Effect of Masked Token Prediction in EEG Tokenization
-We conducted an ablation study on TFM-Encoder to assess the impact of masked token prediction pretraining in a fully discretized framework.
+CHB-MIT Single BIOT 0.6582 $\pm$ 0.0896 0.3127 $\pm$ 0.0890 0.8456 $\pm$ 0.0333
+BIOT-TFM 0.5893 $\pm$ 0.0197 0.3341 $\pm$ 0.0349 $\uparrow$ 0.8752 $\pm$ 0.0123 $\uparrow$
+LaBraM 0.5035 $\pm$ 0.0078 0.0959 $\pm$ 0.0742 0.6624 $\pm$ 0.1050
+LaBraM-TFM 0.5473 $\pm$ 0.047 $\uparrow$ 0.2376 $\pm$ 0.0461 $\uparrow$ 0.7863 $\pm$ 0.0438 $\uparrow$
+Multiple BIOT 0.7068 $\pm$ 0.0457 0.3277 $\pm$ 0.0460 0.8761 $\pm$ 0.0284
+BIOT-TFM 0.6197 $\pm$ 0.0085 0.3484 $\pm$ 0.0078 $\uparrow$ 0.8726 $\pm$ 0.0098
+LaBraM 0.5260 $\pm$ 0.0369 0.2138 $\pm$ 0.0523 0.7750 $\pm$ 0.0540
+LaBraM-TFM 0.5579 $\pm$ 0.0394 $\uparrow$ 0.2445 $\pm$ 0.0351 $\uparrow$ 0.7887 $\pm$ 0.0423 $\uparrow$
+Table 9 presents detailed results on integrating TFM-Tokenizer with BIOT and LaBraM.
+Across all metrics and settings, TFM-Tokenizer improves performance in $93\%$ of cases, demonstrating its effectiveness in enhancing existing EEG foundation models.
+C.5 Effect of Masked Token Prediction in EEG Tokenization
+Figure 8: Masked Token Prediction Ablation
+We conducted an ablation study on downstream transformer to assess the impact of masked token prediction pretraining in a fully discretized framework.
 Using a pretrained TFM-Tokenizer, we compared two approaches: (1) masked token prediction pretraining followed by fine-tuning and (2) direct fine-tuning without pretraining.
@@
 While Cohen’s Kappa and Weighted F1 showed no significant differences between the two approaches, masked token prediction pretraining significantly improved balanced accuracy across all TFM-Tokenizer variants.
-This suggests that pretraining enhances class-wise prediction consistency by capturing token dependencies and making TFM-Encoder more robust to missing channels or time segments, a common challenge in EEG analysis.
-Figure 6. Masked Token Prediction Ablation
-A.3. Removing Position Embedding in TFM-Tokenizer Improves Token Learning
-Table 6. TFM-Tokenizer Comparison with and w/o Position Embedding (PE) on TUEV Dataset
+This suggests that pretraining enhances class-wise prediction consistency by capturing token dependencies and making downstream transformer more robust to missing channels or time segments, a common challenge in EEG analysis.
+C.6 Removing Position Embedding in TFM-Tokenizer Improves Token Learning
+Table 10: TFM-Tokenizer Comparison with and w/o Position Embedding (PE) on TUEV Dataset
 Method Utilization Uniqueness Balanced Cohen’s Weighted
-[…]
+$\%$ (GM) $\%$ Acc. Kappa F1
 TFM-Tokenizer + PE $12.87$ $1.94$ $0.4765\pm 0.038$ $0.5119\pm 0.022$ $0.7457\pm 0.012$
-TFM-Tokenizer w/o PE $9.78$ $2.14$ 0.4943 ±$\pm 0.052$ 0.5337 ±$\pm 0.031$ 0.7570 ±$\pm 0.016$
-Through our empirical analysis, we found that the performance of our TFM-Token significantly improved when no position embedding was applied to the TFM-Tokenizer.
+TFM-Tokenizer w/o PE $9.78$ $2.14$ 0.4943 $\pm 0.052$ 0.5337 $\pm 0.031$ 0.7570 $\pm 0.016$
+Through our empirical analysis, we found that the performance significantly improved when no position embedding was applied to the TFM-Tokenizer.
 EEG patterns are inherently chaotic and non-stationary, meaning similar motifs can occur at any position within the signal.
@@
 This improvement is further illustrated in the motifs captured by the TFM-Tokenizer’s tokens in Figure 5 in Section 4.7.
-A.4. Ablation on Token Vocabulary Size
+C.7 Downstream Model Ablation
+Table 11: Ablation on number of transformer layers in the downstream model
+# Number of Performance Metrics
+Layers Params. Balanced Acc. Cohen’s Kappa Weighted F1
+1 0.58M 0.4486 $\pm$ 0.0297 0.5404 $\pm$ 0.0168 0.7603 $\pm$ 0.0096
+2 0.63M 0.4920 $\pm$ 0.0595 0.5758 $\pm$ 0.0169 0.7780 $\pm$ 0.0089
+4 0.72M 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163
+6 0.82M 0.5025 $\pm$ 0.0592 0.4996 $\pm$ 0.0208 0.7410 $\pm$ 0.0104
+12 1.12M 0.5016 $\pm$ 0.0730 0.5088 $\pm$ 0.0272 0.7456 $\pm$ 0.0139
+We ablated the number of transformer layers in the downstream model on the TUEV dataset, with results presented in Table 11.
+Notably, even with significantly fewer parameters (two layers), the model maintains competitive and, in some cases, better performance across key metrics.
+This highlights the potential for developing lightweight and efficient models for EEG analysis without substantial performance trade-offs.
+C.8 Ablation on Token Vocabulary Size
 To evaluate the impact of token vocabulary size on performance and token learning, we conducted an ablation study by varying the vocabulary size from 256 to 8192 in powers of two.
@@
 Notably, Figure 10b shows that class-token uniqueness scores increase with vocabulary size, contributing to the improvement in balanced accuracy by enabling learning more unique class-specific tokens.
-Figure 7. Token vocabulary size ablation with performance metrics
-Figure 8. Token vocabulary size ablation with token utilization and uniqueness
-A.5. Ablation on Masking
-Table 7. Ablation on masking used for the pretraining of TFM-Tokenizer on TUEV Dataset
+Figure 9: Token vocabulary size ablation with performance metrics
+Figure 10: Token vocabulary size ablation with token utilization and uniqueness
+C.9 Ablation on Masking
+Table 12: Ablation on masking used for the pretraining of TFM-Tokenizer on TUEV Dataset
 Masking Strategy Balanced Acc. Cohen’s Kappa Weighted F1
@@
 However, temporal masking results in a decline in Cohen’s Kappa and Weighted F1, which is then resolved by introducing symmetric masking, achieving the overall best performance.
-Appendix B TFM-Token Implementation and Hyperparameter Tuning
-Figure 9. TFM-Token Overview
-Figure 9 presents an overview of TFM-Token during inference. This section provides additional details on the implementation and training of the framework.
-B.1. Training Pipeline:
-For all experiments, we follow a single-dataset setting, where all processes in each experiment are conducted within the same dataset.
-The training process of our framework is as follows: (1) TFM-Tokenizer unsupervised pretraining, (2) TFM-Encoder pretraining using masked token prediction, and finally (3) fine-tuning on the same dataset for downstream tasks.
-B.2. Hyperparameter Tuning of TFM-Token
-We employed a systematic approach to optimize the hyperparameters of both the TFM-Tokenizer and TFM-Encoder models using Ray Tune [^1] https://docs.ray.io/en/latest/tune/ with the Optuna [^2] https://optuna.org/ search algorithm.
+Appendix D TFM-Tokenizer Implementation and Hyperparameter Tuning
+Figure 11: TFM-Tokenizer framework Overview
+Figure 11 presents an overview of the framework during inference. This section provides additional details on the implementation and training of the framework.
+D.1 Hyperparameter Tuning of TFM-Tokenizer and Downstream Transformer
+We employed a systematic approach to optimize the hyperparameters of both the TFM-Tokenizer and downstream transformer models using Ray Tune [^1] https://docs.ray.io/en/latest/tune/ with the Optuna [^2] https://optuna.org/ search algorithm.
 Our optimization process followed a three-phase strategy.
@@
 This was followed by tuning the training optimizer’s parameters, including learning rate and weight decay.
-The second phase focused on the TFM-Encoder optimization for the classification task, where we first tuned its architectural parameters (depth and number of heads), followed by training the optimizer’s parameters while keeping the tokenizer frozen.
-The third phase focused on tuning optimizer parameters for the masked token prediction pretraining of the TFM-Encoder.
+The second phase focused on the downstream transformer optimization for the classification task, where we first tuned its architectural parameters (depth and number of heads), followed by training the optimizer’s parameters while keeping the tokenizer frozen.
+The third phase focused on tuning optimizer parameters for the masked token prediction pretraining of the downstream transformer.
 To ensure a fair comparison with LaBraM’s neural tokenizer, we maintained a vocabulary size of $8,192$ and an embedding dimension of $64$.
 For our ablation studies involving raw signal-only and STFT-only variants, we doubled the embedding dimensions of the temporal encoder and frequency patch encoder to match the codebook dimension while maintaining all other parameters same.
-Detailed hyperparameter configurations for both TFM-Tokenizer and TFM-Encoder are provided in Appendices B.3 and B.4, respectively.
-B.3. TFM-Tokenizer Hyperparameters
-Table 8. Hyperparameters for TFM-Tokenizer unsupervised pretraining on single-channel setting
+Detailed hyperparameter configurations for both TFM-Tokenizer and downstream transformer are provided in Appendices D.2 and D.3, respectively.
+D.2 TFM-Tokenizer Hyperparameters
+Table 13: Hyperparameters for TFM-Tokenizer unsupervised pretraining on single-channel setting
 Hyperparameter Values
@@
 # of Pretraining Epochs 100
-Table 9. Hyperparameters for TFM-Tokenizer
+Table 14: Hyperparameters for TFM-Tokenizer
 Hyperparameter Values
@@
 Linear Decoder 100
-B.4. TFM-Encoder Hyperparameters
-Table 10. Hyperparameters for TFM-Encoder, its masked token prediction pretraining and downstream finetuning
+D.3 Downstream Transformer Encoder Hyperparameters
+Table 15: Hyperparameters for downstream transformer, its masked token prediction pretraining and downstream finetuning
 Hyperparameter Values
@@
 label smoothing (multi-class) 0.1
-Appendix C Additional Experiment Details
-C.1. Dataset Statistics and Splits
-Table 11. Dataset Summary
-Dataset # of Recordings # of Samples Duration (s) Task
-TUEV $11,914$ $112,491$ $5$ EEG Event Classification
-IIIC Seizure $2,689$ $135,096$ $10$ Seizure Type Classification
-CHB-MIT $686$ $326,993$ $10$ Seizure Detection
-TUAB $2,339$ $409,455$ $10$ Abnoral EEG Detection
-This section provides detailed information on the datasets used in our experiments and their respective splits.
-Table 11 summarizes key statistics, including the number of recordings, the total number of samples after preprocessing, their duration, and the corresponding downstream tasks.
-For TUEV and TUAB, we utilized the official training and test splits provided by the dataset and further divided the training splits into $80\%$ training and $20\%$ validation sets.
-We performed a subject-wise split into $60\%$ training, $20\%$ validation, and $20\%$ test on the IIIC Seizure dataset.
-In the CHB-MIT dataset, we used 1-19 subjects for training, 20-21 for validation, and 22-23 for testing.
-C.2. STFT parameters
-To extract frequency-domain representations of the EEG, we utilized the STFT function from PyTorch.
-The recommendations of (Yang et al., 2024) guided our parameter selection and empirical analysis of different configurations to optimize the time-frequency resolution tradeoff.
-The final parameters are as follows:
-Table 12. STFT parameters
-Parameter Value Description
-FFT size ($n_{\text{fft}},L$) $200$ Number of frequency bins (equal to resampling rate)
-Hop length $H$ $100$ Step size for sliding window ($50\%$ overlap)
-Window type Hann A smoothing window function to reduce spectral leakage
-Output representation Magnitude Only the absolute values of the STFT are retained
-Centering False The STFT is computed without implicit zero-padding
-One-sided output True Only the positive frequency components are kept
+Appendix E More Related Works
+Frequency Representation Collapse.
+Frequency domain analysis is crucial in EEG and general time series analysis  (Elvander & Jakobsson, 2020; Wu et al., 2021; 2023; Woo et al., 2022).
+In real-world signals, time-domain observations inherently mix multiple frequency components, and high-energy, low-frequency signals often dominate the spectrum  (Huang Norden E Shen Zheng & H, 1998; Lai et al., 2018).
+As a result, these entangled frequency features makes it difficult for models to distinguish between them  (Zhou et al., 2022; Piao et al., 2024).
+Recent studies have shown that these entangled signals can lead to a collapse in the learned frequency representations  (Zhi-Qin John Xu et al., 2020; Piao et al., 2024).
+Models tend to overemphasize the dominant low-frequency features while neglecting the high-frequency details.
+This issue can lead to a lack of capturing various EEG waveforms and degenerating data representation  (Park & Kim, 2022).
+Motivated by these works, our paper focuses on developing methods to learn diverse, informative frequency features.
+In Section 7, we provide an analysis of our proposed frequency-domain tokenizer and its impact on model performance.
+Appendix F LLM Usage Statement
+We used large language models (LLMs) solely for writing support, including grammar correction, sentence refinement, and clarity improvements.
+All conceptual contributions, algorithm design, code development, experiments, and analyses were conducted entirely by the authors.